### v1

# An operational approach to classifying measurement incompatibility

Abstract
Measurement incompatibility has proved to be an important resource for information processing tasks. In this work, we analyze various levels of incompatibility of measurement sets. We provide operational classification of measurement incompatibility with respect to two elementary classical operations, viz., coarse-graining of measurement outcomes and convex mixing of different measurements. We derive analytical criteria for determining when a set of projective measurements is fully incompatible with respect to coarse-graining or convex mixing. Robustness against white noise is investigated for mutually unbiased bases that can sustain full incompatibility. Furthermore, we propose operational witnesses for different levels of incompatibility subject to classical operations, using the input-output statistics of Bell-type experiments as well as experiments in the prepare-and-measure scenario.
I introduction
Measurement incompatibility is a concept relating observables which cannot be measured jointly with arbitrary accuracy Gühne et al. (2023). It is a purely quantum effect of which the most well-known example concerns the position and momentum of a quantum particle. Being a fundamental concept of quantum theory, it takes a pivotal role in explaining several quantum phenomena, such as, Bell-nonlocality Fine (1982a); Wolf et al. (2009), Einstein–Podolsky–Rosen steering Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015), measurement uncertainty relations Pramanik et al. (2013); Busch et al. (2014); Maity et al. (2017); Saha et al. (2020), quantum contextuality Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), quantum violation of macrorealism Mal and Majumdar (2016); Uola et al. (2019), and temporal and channel steering Karthik et al. (2015); Banik et al. (2015); Uola et al. (2018). Measurement incompatibility finds applications in various quantum information processing tasks. Recently, it has been shown that measurement incompatibility is necessary for quantum advantage in any one-way communication task Saha et al. (2023).
The significance of measurement incompatibility in various quantum effects calls for its in-depth characterization. Towards this direction, a classification of measurement incompatibility with respect to projection onto subspaces has been recently performed Uola et al. (2021). In the present work, our aim is to classify measurement incompatibility with respect to certain basic classical operations. From an operational perspective, here we employ two such operations, namely coarse-graining of measurement outcomes and convex mixing of measurement settings. Coarse-graining of measurement outcomes arises naturally Toscano et al. (2018) in several instances, such as, for example, quite obviously in measurements on continuous variable system. Though the eigen spectra of the observables are infinite-dimensional taking continuous values, real-world experimental devices are limited by finite precision, leading to the measurement outcomes taking a finite number of discrete values. This inaccuracy in the recording of measurement outcomes is manifested in the coarse-graining of measurement outcomes, which is inevitable in practice. Coarse-graining has also been employed to study the phenomenon of quantum to classical transition. It is observed that quantum phenomena may disappear due to imprecision of measurement outcomes Kofler and Brukner (2007); Rudnicki et al. (2012); Jeong et al. (2014); Mal and Majumdar (2016).
Device imperfection may also lead to the measurement device performing probabilistically a set of measurements, instead of always performing the desired particular measurement. In such a case, a convex mixing of the given set of measurements arises effectively. As incompatibility of measurements is a quantum concept, it is interesting to examine how this property behaves under such elementary classical operations. Our study is motivated by the question as to how one may compare the degree of incompatibility between two different sets of measurements. For instance, if the first set remains incompatible for every possible non-trivial coarse-graining of the measurement outcomes, but the second set becomes compatible for a certain coarse-graining, it follows that the first set of measurements exhibits stronger incompatibility compared to the second one. A similar argument holds for the case of convex-mixing of measurements too.
As incompatible measurements are useful for performing various information processing tasks, given a measurement device claiming to produce incompatible measurements, we must certify it before using it in an information processing task. Verification of incompatiblity of the measurements is possible from the input-output measurement statistics obtained from the device without knowing its internal functioning. Device-independent protocols are conceptually most powerful, relying only on the input-output statistics Brunner et al. (2014), with a wide range of applications from quantum cryptography Ekert (1991); Acín et al. (2007) to communication complexity Buhrman et al. (2010), self-testing Mayers and Yao (2004); Šupić and Bowles (2020); Goswami et al. (2018); Bian et al. (2020) and incompatibility witnesses Quintino et al. (2019). However, they require shared entanglement, an expensive resource, and prohibition of communication between the involved parties. On the other hand, semi-device-independent protocols are inspired by the standard prepare-and-measure scenario, with an additional constraint on the dimension of the communicated quantum states Van Himbeeck et al. (2017); Moreno et al. (2021). They have been applied in quantum key distribution Pawłowski and Brunner (2011), quantum networks Bowles et al. (2015), randomness certification Passaro et al. (2015), random access codes Li et al. (2012), and certification of meassurements Maity et al. (2021); Das et al. (2022).
In the present work we investigate the issue of certification of measurement incompatibility subjected to the classical operations of coarse-graining and convex mixing of measurements from both the device-independent and the semi-device-independent perspective. An essential ingredient in the performance of quantum information protocols is the presence of noise. Within the context of our present study, noise is reflected in the degrading of incompatibility properties of various measurement sets. We examine the critical amount of robustness of the noisy measurement above which the measurements remain incompatible under the mentioned classical operations. We then study incompatibility witnesses in the device-independent and the semi-device-independent scenarios.
Our paper is structured as follows. In Sec. (II) we present the basic mathematical ingredients for classifying measurement incompatibility under the two operations of coarse-graining of outcomes and convex-mixing of measurements, respectively. In Sec. (III) we study the robustness under noise for various levels of incompatibility subjected to the above classical operations. In Sec. (IV) we define operational witnesses of incompatibility, and furnish examples to study their performance in both device-independent and the semi-device-independent frameworks. Concluding remarks are presented in Sec. (V).
II Classifying measurement incompatibility under elementary classical operations
In the case of projective measurements, the observables whose corresponding operators commute are jointly measurable. However, this is not true in the case of general quantum measurements. The most general quantum measurement is described by Positive Operator Valued Measures (POVM), which is a set of operators, $\{O_{\alpha}\}$ with $0\leq O_{\alpha}\leq\mathbbm{1}$, and $\sum_{\alpha}O_{\alpha}=\mathbbm{1}$. A set of measurements defined by $\{M_{z_{x}|x}\}$ (here $x$ corresponds to the measurement, $z_{x}$ denotes the outcomes of the measurement labeled by $x$) is compatible if there exists a parent POVM, $G_{\lambda}$ and classical post-processing for each $x$ given by $\{p(z_{x}|x,\lambda)\}$ such that
$\forall z_{x},x,\quad M_{z_{x}|x}=\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda},$ (1)
where $0\leq p(z_{x}|x,\lambda)\leq 1,\sum_{z_{x}}p(z_{x}|x,\lambda)=1\forall x,\lambda$    Gühne et al. (2023).
II.1 Coarse-graining of outcomes
Coarse-graining of the outcomes of a measurement is a concept where some of the outcomes are treated on the same footing, i.e., different outcomes are clubbed together to form a single outcome. After the coarse-graining of the outcomes, the effective number of outcomes reduces. Coarse-graining may occur in practice due to measurement errors. Additionally, coarse-graining is used in when one does not require a finer description of the values of the measurement outcomes, but rather a broad description suffices to understand the nature of some physical property.
Let us introduce the notation $[k]:=\{0,\cdots,k-1\}$ for any natural number $k.$ Consider a $\overline{d}$ outcome measurement $\{M_{\overline{z}}\}$ and an arbitrary coarse-graining that yields a $d$ outcome measurement $\{M_{z}\}$, where $\overline{z}\in[\overline{d}]\,\,\text{and}\,\,z\in[d]$. The coarse-graining can be defined by conditional probabilities $\{c(z|\overline{z})\}$ such that
$M_{z}=\sum_{\overline{z}}c(z|\overline{z})M_{\overline{z}},$ (2)
and $c(z|\overline{z})\in\{0,1\}$. Trivial coarse-graining is the one where the measurement becomes single outcome, that is, $c(z|\overline{z})=1,\forall\overline{z}$. Two incompatible measurements may not necessarily remain incompatible after certain coarse-graining of their outcomes.
Definition 1 (Fully incompatible measurements w.r.t coarse-graining).
A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is fully incompatible with respect to (w.r.t) coarse-graining if they remain incompatible after all possible nontrivial coarse-graining. That is, if the resultant set of measurements given by
$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},$ (3)
after all possible sets of nontrivial coarse-graining $\{c_{x}(z_{x}|\overline{z}_{x})\}$ pertaining to every setting $x$ is incompatible, then we call them fully incompatible.
Definition 2 ($d$-incompatible measurements w.r.t coarse-graining).
A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is $d$-incompatible w.r.t coarse-graining if they remain incompatible after all possible nontrivial coarse-graining that give rise to at least $d$ outcome measurements. In other words, if the resultant set of measurements
$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},\ z_{x}\in[k_{x}],k_{x}\geqslant d,\forall x$ (4)
after all possible coarse-graining resulting in at least $d$ outcome measurements is incompatible, then we call them $d$-incompatible.
For example, the four outcome rank-one projective measurement pair, defined by the following (un-normalized) vectors
$M=\Big{\{}|0\rangle,|1\rangle,|2\rangle,|3\rangle\Big{\}}$ (5)
and
$N=\Big{\{}|0+1\rangle,|0-1\rangle,|2+3\rangle,|2-3\rangle\Big{\}}$ (6)
is 3-incompatible, where the notation $|0+1\rangle=|0\rangle+|1\rangle,$ and similarly for others.
Observation 1.
A set of fully incompatible measurements is equivalent to 2-incompatible measurements w.r.t coarse-graining.
If a set of measurements is 2-incompatible, then it implies that the set remains incompatible after all possible coarse-graining of the outcomes such that the number of outcomes of each measurement in the newly formed set of measurements is greater than or equal to two. Also, the lowest number of outcomes of measurement is two for a nontrivial coarse-graining. Furthermore, if a set of measurements is $d$-incompatible, then, by definition, it is $n$-incompatible as well, where $n>d$, but the reverse is not true. This proves Observation 1.
Observation 2.
[…]
$\left[\sum_{i\in\mathcal{M}_{k}}P_{i},\sum_{j\in\mathcal{N}_{l}}Q_{j}\right]\neq 0,\ \forall k,l.$ (7)
Proof.
[…]
Theorem 1.
The following condition is necessary but not sufficient for two rank-one projective measurements, defined by $\{|\psi_{i}\rangle\}$ and $\{|\phi_{j}\rangle\}$, to be fully incompatible w.r.t coarse-graining:
$\langle\psi_{i}|\phi_{j}\rangle\neq 0,\ \forall i,j.$ (8)
The above condition is necessary and sufficient for two 3-dimensional rank-one projective measurements.
Proof.
First, note that for sharp measurement compatibility and commutativity are equivalent Heinosaari and Wolf (2010). Suppose $\exists i,j,$ such that $\langle\psi_{i}|\phi_{j}\rangle=0$. Consider coarse-graining of all other outcomes except $i$ and $j$ for the two measurements. Then, the resulting measurements will be $\{|\psi_{i}\rangle\!\langle\psi_{i}|,\mathbbm{1}-|\psi_{i}\rangle\!\langle\psi_{i}|\}$ and $\{|\phi_{j}\rangle\!\langle\phi_{j}|,\mathbbm{1}-|\phi_{j}\rangle\!\langle\phi_{j}|\}$, which commute with each other (i.e., the resulting measurements are compatible) since $\langle\psi_{i}|\phi_{j}\rangle=0$. Thus, if the measurements are fully incompatible, condition (8) holds.
To show that (8) is not sufficient, consider the following two $4$-dimensional rank-one projective measurements (with the normalization factor $1/\sqrt{2}$),
$\{|0\rangle+|1\rangle,|0\rangle-|1\rangle,|2\rangle+|3\rangle,|2\rangle-|3\rangle\}$
$\{|0\rangle+|2\rangle,|0\rangle-|2\rangle,|1\rangle+|3\rangle,|1\rangle-|3\rangle\}.$ (9)
One can check that (8) holds for all $i,j=0,1,2,3$. But coarse-graining of outcomes $0,1$ and $2,3$ for both the measurements leads to
$\{|0\rangle\!\langle 0|+|1\rangle\!\langle 1|,|2\rangle\!\langle 2|+|3\rangle\!\langle 3|\},$
$\{|0\rangle\!\langle 0|+|2\rangle\!\langle 2|,|1\rangle\!\langle 1|+|3\rangle\!\langle 3|\},$ (10)
which are compatible.
[…]
II.2 Convex mixing of measurements
The concept of convex mixing of measurements may be best understood by considering an example: suppose we have a measurement device that produces three different measurements $(M_{0},M_{1}\,\,\text{and}\,\,M_{2})$ and has two input switches. If the first switch is ON it performs measurement $M_{0}$ with probability $q$ and $M_{1}$ with probability $(1-q)$ and for the second switch, it always performs $M_{2}.$ Thus, in principle, the device produces two measurements: $M_{2}$ and a convex mixture of $M_{0}$ and $M_{1}.$
Suppose $M_{0}=\{M_{z|0}\},M_{1}=\{M_{z|1}\},M_{2}=\{M_{z|2}\}$ such that they have same number of outcomes $z\in[d]$. Now the measurement that is realized by a convex combination of $M_{0}$ and $M_{1}$ with some weightage $q\,\,\text{and}\,\,1-q$, where $q\in[0,1]$is defined as
$M_{(0,1,\pi)}=\{qM_{z|0}+(1-q)M_{\pi(z)|1}\},$ (11)
where $\pi(z)$ is any permutation on outcome $z$. Here $\pi(z)$ is a bijective function from the set $z\in[d]$ to itself. Even if $M_{2}$ is incompatible with $M_{0}$ and $M_{1}$ separately, $M_{2}$ is not necessarily incompatible with $M_{(0,1,\pi)}$ for all values of $q$.
Definition 3.
[…]
Consider three unbiased qubit measurements,
$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\vec{n_{i}}\cdot\vec{\sigma}\right),\quad i=0,1,2,$ (12)
and $||\vec{n_{i}}||\leqslant 1$. These three measurements are fully incompatible w.r.t convex mixtures if and only if (see eq.(7) of Gühne et al. (2023)),
$||\vec{n}_{i}+q\vec{n}_{j}\pm(1-q)\vec{n}_{k}||+||\vec{n}_{i}-q\vec{n}_{j}\mp(1-q)\vec{n}_{k}||>2,$ (13)
for all $q$, and for all $(i,j,k)\in\{(0,1,2),(1,2,0),(2,0,1)\}$.
Observation 3.
If three qubit measurements (12) are such that $\vec{n}_{i}$ are in the same plane of the Bloch sphere, then they are not fully incompatible w.r.t convex mixtures.
Proof.
Note that permutation of the outcomes of $M_{i}$ in (12) is nothing but taking $-\vec{n}_{i}$ instead of $\vec{n}_{i}$. If the three $\vec{n}_{i}$ are in the same plane, then there exists at least one triple $(i,j,k)$ and permutations of outcomes such that
$\vec{n}_{0}=\frac{1}{c}(q\vec{n}_{1}+(1-q)\vec{n}_{2})$ (14)
for some $c\in(0,1],q\in[0,1]$. In other words, there exists a triple $(i,j,k)$ and permutation of outcomes such that $\vec{n}_{i}$ is expressed as a linear combination of $\vec{n}_{j}$ and $\vec{n}_{k}$ with non-negative coefficients (here $q/c$ and $(1-q)/c$), such that sum of those two non-negative coefficients is greater than or equal to 1. Substituting this into left hand side of (13) taking $+$ sign, we find
$||(1+c)\vec{n}_{i}||+||(1-c)\vec{n}_{i}||\leqslant|1+c|+|1-c|=2.$ (15)
[…]
Observation 4.
Suppose three qubit measurements (12) are such that $\vec{n}_{0}=\nu_{0}\hat{x}$, $\vec{n}_{1}=\nu_{1}\hat{y}$, $\vec{n}_{2}=\nu_{2}\hat{z}$ with $0\leq\nu_{1},\nu_{2},\nu_{3}\leq 1$, that is, they are the noisy version of Pauli observables,
$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\nu_{i}\sigma_{i}\right)=\nu_{i}\left(\frac{\mathbbm{1}\pm\sigma_{i}}{2}\right)+(1-\nu_{i})\frac{\mathbbm{1}}{2},$ (16)
taking $i=0,1,2$ and $(\sigma_{0},\sigma_{1},\sigma_{2})=(\sigma_{x},\sigma_{y},\sigma_{z})$. Then they are fully incompatible w.r.t convex mixture if and only if
$\min\bigg{\{}\nu^{2}_{0}+\frac{\nu^{2}_{1}\nu^{2}_{2}}{\nu^{2}_{1}+\nu^{2}_{2}},\nu^{2}_{1}+\frac{\nu^{2}_{0}\nu^{2}_{2}}{\nu^{2}_{0}+\nu^{2}_{2}},\nu^{2}_{2}+\frac{\nu^{2}_{0}\nu^{2}_{1}}{\nu^{2}_{0}+\nu^{2}_{1}}\bigg{\}}>1.$ (17)
Proof.
In terms of $\nu_{i}$, the left hand side of (13) becomes $2\sqrt{\nu^{2}_{i}+q^{2}\nu^{2}_{j}+(1-q)^{2}\nu^{2}_{k}}$. Note that the minimum of $q^{2}\nu^{2}_{j}+(1-q)^{2}\nu^{2}_{k}$ occurs at $\tilde{q}=\nu_{k}^{2}/(\nu^{2}_{j}+\nu^{2}_{k})$. Since $0\leq\nu_{k}^{2}/(\nu^{2}_{j}+\nu^{2}_{k})\leq 1$ for any $\nu_{j},\nu_{k}\in[0,1]$, the above-mentioned minimum can always be achieved with a suitable choice of $q$. Hence, we have that
$2\sqrt{\nu^{2}_{i}+q^{2}\nu^{2}_{j}+(1-q)^{2}\nu^{2}_{k}}$ $\geq 2\sqrt{\nu^{2}_{i}+\tilde{q}^{2}\nu^{2}_{j}+(1-\tilde{q})^{2}\nu^{2}_{k}}$
$=2\sqrt{\nu^{2}_{i}+\frac{\nu^{2}_{j}\nu^{2}_{k}}{\nu^{2}_{j}+\nu^{2}_{k}}}$ (18)
Thus, the r.h.s of (13) is greater than 2 for all values of $q$ if and only if
$\nu^{2}_{i}+\frac{\nu^{2}_{j}\nu^{2}_{k}}{\nu^{2}_{j}+\nu^{2}_{k}}>1.$ (19)
[…]
Clearly, the three Pauli observables are fully incompatible w.r.t. convex mixture, and moreover, if we take an equal amount of noise $\nu_{0}=\nu_{1}=\nu_{2}=\nu$, then (17) implies $\nu>\sqrt{2/3}$.
We can generalize this notion for a set of $n$ measurements.
Definition 4 ($k$-incompatible measurements w.r.t. convex mixtures).
A set of measurements is $k$-incompatible w.r.t convex mixtures if after taking all possible convex mixtures over all possible $k$ partitions the resulting $k$ number of measurements from the set of $n$ measurements (with relabelling of outcomes) is incompatible.
Definition 5 (Fully incompatible measurements w.r.t. convex mixtures).
A set of measurements is fully incompatible w.r.t convex mixtures if it remains incompatible after taking all possible convex mixtures over all possible partitions of $n$ measurements (with relabelling of outcomes).
Theorem 2.
Fully incompatible measurements w.r.t convex mixtures imply that every pair of measurements from that set is incompatible. The reverse implication does not hold.
Consider a set of $n$ measurements, $M=\{M_{1},M_{2},M_{3},\ldots,M_{n}\}$, in which there is a pair of measurements $M_{i}$ and $M_{j}$ which are compatible. Now if we make two partitions where $M_{i}$ and $M_{j}$ are in different partitions, the convex mixture of measurements must be compatible. This is true for any compatible pair $M_{i}$ and $M_{j}$. Thus, if the measurements are fully incompatible w.r.t convex mixture, every pair of measurements must necessarily be incompatible.
The reverse is not true. Consider the three noisy Pauli measurements of Eq.(16). It can be shown by using semi-definite programming that if $0.71<\nu\leq 0.81$, the measurements are pairwise incompatible but it is not fully incompatible w.r.t convex mixture.
III Robustness under noise for different levels of incompatibility
In this section, we analyze the role of noise on quantum measurements and study how the incompatibility properties depend on it. Due to the ubiquitous nature of noise, it is pertinent to study the extent to which noise could be tolerated by a set of measurements while still retaining their incompatibility. We take a noisy version of mutually unbiased bases measurements, $M_{i|x}\equiv\{\nu|\phi_{i}^{x}\rangle\langle\phi_{i}^{x}|+\frac{(1-\nu)}{d}\mathbbm{1}_{d}\},\,\,\text{where}\,\,\{|\phi_{i}^{x}\rangle\}$ form mutually unbiased bases measurements in $\mathbb{C}^{d}$. Here $\nu$ is the robustness parameter (or visibility parameter) and $(1-\nu)$ is the noise parameter, $0\leq\nu\leq 1$. When the noise parameter is zero (i.e., robustness, $\nu=1$), the measurements are fully incompatible, and when the noise parameter is one (i.e., the robustness, $\nu=0$), the measurements are trivial and compatible. Our aim is to obtain the critical value of the robustness parameter above which the measurements remain incompatible.
To check the compatibility, i.e., the existence of a parent POVM, we use the method described in Gühne et al. (2023). This can be casted as a semi-definite programming (SDP) problem that takes a set of measurements $\{M_{z_{x}|x}\}$ and deterministic classical post-processings $p(z_{x}|x,\lambda)$ as input, and checks whether the measurements are compatible or not, subject to the constraints
$\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda}=M_{z_{x}|x}\forall x,z_{x},$ (20)
$\sum_{\lambda}G_{\lambda}=\mathbbm{1},$ (21)
$G_{\lambda}\geq\mu\mathbbm{1},$ (22)
where $\mu$ is the optimization parameter. This method finds the maximum value of $\mu$ for each $\{p(z_{x}|x,\lambda)\}$. If this optimization returns a negative value of $\mu$, then the constraint of Eq.(22) can’t be fulfilled, which implies that the measurements $\{M_{z_{x}|x,\lambda}\}$ are incompatible. Otherwise, they are compatible.
III.1 Coarse-graining of outcomes
Dimension $3$. Let $\{P_{i}\}$ and $\{Q_{j}\},i,j\in[3]$ be two measurements acting on $\mathbb{C}^{3}$, where
$P_{i}=\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{3},$
$Q_{j}=\nu|\psi_{j}\rangle\langle\psi_{j}|+(1-\nu)\frac{\mathbbm{1}}{3},$ (23)
with
$|\psi_{0}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+|1\rangle+|2\rangle),$
$|\psi_{1}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega|1\rangle+\omega^{2}|2\rangle),$
$|\psi_{2}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega^{2}|1\rangle+\omega|2\rangle),$
It is observed that for $\nu>0.683$, the two measurements are incompatible. For finding ”$2$-incompatible w.r.t coarse-graining”, we consider all possible coarse-graining of outcomes of the two measurements such that after coarse-graining, the new set of measurements have at least two outcomes each. We find that these measurements are incompatible w.r.t all possible coarse-graining for $\nu>0.711$. Note that for $0.683<\nu\leq 0.711$, the incompatible measurements become compatible under certain coarse-graining of outcomes (see TABLE 1).
Dimension $4$. For checking incompatibility w.rt coarse-graining in $\mathbb{C}^{4}$, the same procedure is repeated for two POVM measurements with $4$ outcomes each. The corresponding measurements are $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\},i,j\in[4]$.
$P^{\prime}_{i}$ $=$ $\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{4},$
$Q^{\prime}_{j}$ $=$ $\nu|\psi^{\prime}_{j}\rangle\langle\psi^{\prime}_{j}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (24)
with
$|\psi^{\prime}_{0}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle+|2\rangle+|3\rangle),$
$|\psi^{\prime}_{1}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle-|2\rangle-|3\rangle),$
$|\psi^{\prime}_{2}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle-|2\rangle+|3\rangle),$
$|\psi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+|2\rangle-|3\rangle),$ (25)
for $\nu>0.666$, these measurements are incompatible. We find that for $\nu>0.675$, the measurements are incompatible under all possible non-trivial coarse-graining that gives rise to at least $3$ outcomes. The measurements are fully incompatible for $\nu>0.72$. The results are summarized in TABLE 1.
In dimension $3$ In dimension $4$
Incompatible 0.683 0.666
2-incompatible 0.711 0.720
3-incompatible 0.683 0.675
Table 1: Critical values of robustness for MUBs w.r.t coarse-graining in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
III.2 Convex mixing of measurements
Dimension $3$. For checking incompatibility w.r.t convex mixtures (CM), a minimum of three measurements are needed. The two measurements are $\{P_{i}\}$ and $\{Q_{j}\}$, as given in Eqs.(23), The third measurement is $\{R_{k}\},k\in[3]$ where
$R_{k}=\nu|\phi_{k}\rangle\langle\phi_{k}|+(1-\nu)\frac{\mathbbm{1}}{3}.$ (26)
with
$\begin{split}&|\phi_{0}\rangle=\frac{\omega|0\rangle+|1\rangle+|2\rangle}{\sqrt{3}},|\phi_{1}\rangle=\frac{|0\rangle+\omega|1\rangle+|2\rangle}{\sqrt{3}},\\ &|\phi_{2}\rangle=\frac{|0\rangle+|1\rangle+\omega|2\rangle}{\sqrt{3}},\end{split}$ (27)
$\omega$ being the cube roots of unity.
The results are summarized in TABLE 2.
Dimension $4$. Consider the measurements $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\}$ as defined in Eqs.(III.1) and(III.1) and another measurement $\{R^{\prime}_{k}\}$ where $i,j,k\in[4]$.
$\{R^{\prime}_{k}\}=\nu|\phi^{\prime}_{k}\rangle\langle\phi^{\prime}_{k}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (28)
with
$|\phi^{\prime}_{0}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle-\mathbbm{i}|2\rangle-\mathbbm{i}|3\rangle),$
$|\phi^{\prime}_{1}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+\mathbbm{i}|2\rangle+\mathbbm{i}|3\rangle),$
$|\phi^{\prime}_{2}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle+\mathbbm{i}|2\rangle-\mathbbm{i}|3\rangle),$
$|\phi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle-\mathbbm{i}|2\rangle+\mathbbm{i}|3\rangle).$ (29)
The results are summarized in TABLE 2.
In dimension $3$ In dimension $4$
Incompatible 0.537 0.692
2-incompatible 0.764 0.705
3-incompatible 0.537 0.692
Table 2: Critical values of robustness for MUBs w.r.t convex-mixing in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
IV Operational witnesses of incompatibility
IV.1 Coarse-graining of outcomes
As we mentioned earlier, measurements that are fully incompatible w.r.t coarse-graining show stronger incompatibility compared to the measurements that are not fully incompatible w.r.t coarse-graining. To operationally certify whether a set of measurements are fully incompatible w.r.t coarse-graining is important from the perspective of determing the practical utility of such a set for revealing phenomena such as Bell inequality, steering and contextuality Fine (1982a); Wolf et al. (2009); Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015); Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), as well as for checking the proficiency of such a set for information processing tasks Carmeli et al. (2020); Saha et al. (2023). Operational certification means that we need to infer the incompatibility of the measurements from the input-output statistics of the measurement device without knowing its internal functioning. Below we study the two classes of operational certification separately
Device-independent witness
Here one requires neither any prior knowledge of the internal functioning of the measurement device, nor any idea of the dimension of the system on which the measurements act. In Wolf et al. (2009), it has been proven that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality. We can use this result to witness incompatibility for any two 3-outcome projective measurements before and after coarse-graining, from the input-output statistics.
Semi-Device-independent witness
Here, although we do not have any prior knowledge of the internal functioning of the measurement device, we do know the dimension of the system on which the measurements act. Any two incompatible rank-one projective measurements provide an advantage in $(2,d,d)-$RAC Saha et al. (2020). If any two POVMs each with $\overline{d}$ outcomes acting on $\mathbb{C}^{d}$ are jointly measurable, the average maximum success probability in $(2,\overline{d},d)$ RAC is bounded by Saha et al. (2023)
$P(2,\overline{d},d)\leq\frac{1}{2}(1+\frac{d}{\overline{d}^{2}})=P_{CB}(2,\overline{d},d),$ (30)
where $P_{CB}(2,\overline{d},d)$ is the maximum average success probability obtained in a $(2,\overline{d},d)$ RAC using two compatible measurements. For full incompatibility of two 3-outcome rank-one projective measurements, we prove the following:
Theorem 3.
Two rank-one projective measurements, $M=\Big{\{}|\phi_{0}\rangle,|\phi_{1}\rangle,|\phi_{2}\rangle\Big{\}}$ and $N=\Big{\{}|\psi_{0}\rangle,|\psi_{1}\rangle,|\psi_{2}\rangle\Big{\}}$, can be witnessed to be fully incompatible w.r.t. coarse-graining via RAC iff $20<|\langle\phi_{i}|\psi_{j}\rangle|<\frac{4}{5},\quad\forall\,\,i,j=0,1,$ 2 .
Proof.
Without loss of generality, any pair of three outcome rank-one projective measurements can be written up to unitary freedom as,
$M=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|,|2\rangle\langle 2|\Big{\}}\equiv\Big{\{}|\phi_{i}\rangle\langle\phi_{i}|\Big{\}},$
and
$N=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|,|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}\equiv\Big{\{}|\psi_{j}\rangle\langle\psi_{j}|\Big{\}},$ (31)
where $|\psi_{j}\rangle=\alpha_{j}|0\rangle+\beta_{j}|1\rangle+\gamma_{j}|2\rangle$.
Let us first coarse grain the second and third outcomes. So, the new measurement pair becomes:
$M^{(2,3)}=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|+|2\rangle\langle 2|\Big{\}},$ (32)
$N^{(2,3)}=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}.$
The new compatibility bound in this case is,
[…]
and the bound is tight. $M^{(2,3)}\,\,\text{and}\,\,N^{(2,3)}$ will give advantage in the (2,2,3) RAC if:
$\sum_{x,y=0}^{1}$ $||M^{(2,3)}(x)+N^{(2,3)}(y)||>7,\,\,\text{which implies}$
$||A||$ $+$ $||B||+||C||+||D||>7$ (33)
$\text{with}\,\,A$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(0)$
$=$ $|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
Now, $|\psi_{1}\rangle$ can be written as:
$|\psi_{1}\rangle=\langle 0|\psi_{1}\rangle|0\rangle+\langle u|\psi_{1}\rangle|u\rangle\,\,\text{with}$ (34)
$\langle 0|u\rangle=0\,\,\text{and}\,\,|\langle 0|\psi_{1}\rangle|^{2}+|\langle u|\psi_{1}\rangle|^{2}=1.$
$A$ can be written as a $(2\times 2)$ matrix in the basis $\{|0\rangle,|u\rangle\},$
$A=\begin{pmatrix}1+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (37)
So, the maximum eigenvalue of $A\,\,\text{is}\,\,1+|\langle 0|\psi_{1}\rangle|.$ Thus,
$||A||=1+|\langle 0|\psi_{1}\rangle|.$
By Eq.(IV.1) we have
$B$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(1)$ (38)
$=$ $|0\rangle\langle 0|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|$
$=$ $\mathbbm{1}+|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|.$
Similarly, $B$ can be expressed in a block diagonal matrix in the ortho-normal basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$B=\begin{pmatrix}\Gamma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Gamma=\begin{pmatrix}2+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (43)
and
$||B||=2+|\langle 0|\psi_{1}\rangle|.$
Also, from Eq.(IV.1)
$C$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(0)$ (44)
$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{1}\rangle\langle\psi_{1}|\,\,$
$=$ $\mathbbm{1}-|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
Similarly, $C$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$C=\begin{pmatrix}\Sigma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Sigma=\begin{pmatrix}|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (49)
and
$||C||=1+\sqrt{1-|\langle 0|\psi_{1}\rangle|^{2}}.$
Similarly, from Eq.(IV.1)
$D$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(1)$ (50)
$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\,\,$
$=$ $2\mathbbm{1}-|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$
and $D$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$D=\begin{pmatrix}\Xi&0\\ 0&2\end{pmatrix}\,\,\text{where}\,\,\Xi=\begin{pmatrix}1-|\langle 0|\psi_{1}\rangle|^{2}&-\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ -\langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&2-|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (55)
and $||D||=2.$
[…]
Device-independent witness of full-incompatibility of two measurements w.r.t coarse-graining from a single experiment
We now study the following problem: given two measurements of outcomes $d_{1}$ and $d_{2}$ respectively, how can one witness the full-incompatibility w.r.t coarse-graining in a device-independent way from a single experiment? Here ”single experiment” implies the Bell test which gives a Bell-inequality violation for the two measurements of outcomes $d_{1}$ and $d_{2}$ respectively as chosen as inputs of Alice and some measurement choices for Bob and a shared entangled state. As we know if two measurements are $2-$ incompatible w.r.t coarse-graining then they are fully-incompatible w.r.t coarse-graining. Also, we use the result of Wolf et al. (2009), that any two binary outcome incompatible measurements violate the Bell-CHSH inequality. Thus, from the original measurement statistics we do all possible coarse-grainings to make it two outcome measurements and check the Bell-CHSH inequalities. If, all possible coarse-grained two-outcome measurements violate the Bell-CHSH inequalities then the measurements are fully-incompatible w.r.t coarse-graining.
To understand how this technique works, we consider a well-known example of two three-outcome rank-one projective measurements in $\mathbb{C}^{3}$ which give maximum violation of the CGLMP inequality Collins et al. (2002), a well-studied Zohren and Gill (2008); Das et al. (2019); Karczewski et al. (2022) Bell-type inequality. The projective measurements of Alice are: $A_{a}\equiv\{|\xi\rangle_{A,a}\},$ where
$|\xi\rangle_{A,a}=\frac{1}{\sqrt{3}}\sum_{j=0}^{2}\text{exp}\bigg{(}\mathbbm{i}\frac{2\pi}{3}j(\xi+\alpha_{a})\bigg{)}|j\rangle_{A},$ (56)
with $a\in\{1,2\}$ corresponding to two different measurement settings of Alice, and $\xi\in\{0,1,2\},\alpha_{1}=0,\alpha_{2}=\frac{1}{2}$. The projective measurements of Bob are: $B_{b}\equiv\{|\eta\rangle_{B,b}\},$ where
$|\eta\rangle_{B,b}=\frac{1}{\sqrt{3}}\sum_{j=0}^{2}\text{exp}\bigg{(}\mathbbm{i}\frac{2\pi}{3}j(-\eta+\beta_{b})\bigg{)}|j\rangle_{B},$ (57)
with, $b\in\{1,2\}$ corresponding to two different measurement settings of Bob, and $\eta\in\{0,1,2\},\beta_{1}=\frac{1}{4},\beta_{2}=-\frac{1}{4}.$
The incompatibility of the measurements of Eq.(56) before coarse-graining is guaranteed by the CGLMP-inequality violation. Now we investigate the incompatibility status of the measurements of Alice after all possible coarse-graining from the measurement statistics obtained in the CGLMP experiment. Alice’s measurements now have two outcomes each and it is a (2,2,2,3) scenario, i.e., two inputs for Alice and two inputs for Bob and each input of Alice has two outcomes and each input of Bob has three outcomes. For this scenario, Collins and Gisin have shown that there are a total of 72 CH-facet inequalities Collins and Gisin (2004). The CH-inequalities Fine (1982b) are of the form:
$-1\leq$ $P(00|A_{1},B_{1})+P(00|A_{1},B_{2})+P(00|A_{2},B_{2})$ (58)
$-P(00|A_{2},B_{1})-P(0|A_{1})-P(0|B_{2})\leq 0.$
The other CH-inequalities are obtained by (1) interchanging $A_{1}$ with $A_{2}$, (2) interchanging $B_{1}$ with $B_{2}$, and (3) interchanging both $A_{1}$ with $A_{2}$ and $B_{1}$ with $B_{2}$. Consider the scenario where we coarse-grain (0,1) outcomes for both the inputs of Alice. Lets make the following relabeling $(0,1)\equiv\overline{0}$ and $2\equiv\overline{1}$ for the outcomes of $A_{1}$ and $A_{2}$ and also consider the clubbing of $(0,1)\equiv\overline{0}$ outcomes both for $B_{1}$ and $B_{2}.$ Under this relabelling Eq.(58) takes the form:
$-1\leq$ $P(\overline{0}\,\,\overline{0}|A_{1},B_{1})+P(\overline{0}\,\,\overline{0}|A_{1},B_{2})+P(\overline{0}\,\,\overline{0}|A_{2},B_{2})$ (59)
$-P(\overline{0}\,\,\overline{0}|A_{2},B_{1})-P(\overline{0}|A_{1})-P(\overline{0}|B_{2})\leq 0,$
where,
$P(\overline{0}\,\,\overline{0}|A_{i},B_{j})=P(00|A_{i},B_{j})+P(01|A_{i},B_{j})$
$+P(10|A_{i},B_{j})+P(11|A_{i},B_{j}),$ (60)
$P(\overline{0}|A_{i})=P(0|A_{i})+P(1|A_{i}),$
$P(\overline{0}|B_{i})=P(0|B_{i})+P(1|B_{i}),\,\,i,j\in\{1,2\}.$ (61)
Similarly, there are other 8 possible clubbings for Bob’s measurement outcomes and for each clubbing, we have facet inequalities similar to Eq.(59). One can check that when there are the same coarse-graining of outcomes for both measurement inputs of Alice, we get a CH-inequality violation. Thus, under these coarse-grainings, the two measurements of Alice remain incompatible. When there are different coarse-grainings, for some cases we get CH-violation, but for other cases, we do not get CH-violation. CH-violation under a particular coarse-graining signifies that the measurements are incompatible. However, when there is no CH-violation we can not conclude anything regarding the incompatibility status. This is depicted in TABLE   3.
$A_{1}$ $A_{2}$ Incompatible
(0,1) (0,1) yes
(0,1) (1,2) ?
(0,1) (0,2) yes
(1,2) (0,1) yes
(1,2) (1,2) yes
(1,2) (0,2) ?
(0,2) (0,1) ?
(0,2) (1,2) yes
(0,2) (0,2) yes
Table 3: The incompatibility status for all possible coarse-graining (CG) of outcomes for Alice’s measurements ($A_{1}$ and $A_{2}$) as inferred from the statistics of CGLMP experiment. Here “yes” denotes that under the particular CG of $A_{1}$ and $A_{2}$, the measurement remains incompatible. On the other hand, “????” denotes that we can not make any conclusion about their incompatibility.
IV.2 Convex mixing of measurements
Device-independent witness
Here, one can simply use the result that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality Wolf et al. (2009). Considering the convex mixture of measurements for any three binary-outcome measurements, this result can be used to witness full incompatibility w.r.t. convex mixing of measurements.
Semi-device-independent witness
For three or more numbers of measurements we can witness whether the measurements are fully-incompatible w.r.t convex-mixture in a semi-device independent manner by constructing a suitable random access codes task.
Theorem 4.
Three noisy Pauli measurements of Eq.(16) are witnessed to be fully incompatible w.r.t convex-mixing via RAC iff: $\sqrt{\frac{2}{3}}<\nu\leq 1.$
Proof.
Let us denote the measurements of Eq.(16) as $\mathcal{M}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{0}\sigma_{x}\right)$, $\mathcal{M}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{0}\sigma_{x}\right)$, $\mathcal{N}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{1}\sigma_{y}\right)$, $\mathcal{N}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{1}\sigma_{y}\right)$, $\mathcal{R}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{2}\sigma_{z}\right)$, and $\mathcal{R}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{2}\sigma_{z}\right)$, and consider the measurements $Q^{i}$ with $i\in\{1,2\}$, to be formed by the convex-mixing of $\mathcal{M}$ and $\mathcal{N}$ as
$Q^{1}=\{p\,\,M_{0}+(1-p)\,\,N_{0},\,\,pM_{1}+(1-p)N_{1}\},$
$Q^{2}=\{p\,\,M_{0}+(1-p)N_{1},p\,\,M_{1}+(1-p)N_{0}\}.$
In order to witness the incompatibility status of $(Q^{i},R)$, one can construct a (2,2,2) RAC task. Now, when Bob has to guess Alice’s 1st bit, he performs the measurement $Q^{i}$ which can be realized by the convex mixing of measurements $\mathcal{M}$ and $\mathcal{N}$ as follows: Bob tosses a biased coin which gives a ‘Head’ with probability $p$ and a ‘Tail’ with probability $(1-p)$. On obtaining ‘Head’, Bob performs measurement $\mathcal{M}$, and for ‘Tail’ he performs measurement $\mathcal{N}$. For the second bit, he performs measurement $\mathcal{R}$. The maximum average success probability $P(2,2,2)$ that can be obtained in the RAC task (considering $\nu_{0}=\nu_{1}=\nu_{2}$) is:
$P(2,2,2)$ $=$ $\frac{1}{8}\sum_{j,k=0}^{1}||Q^{i}_{j}+R_{k}||$ (62)
$=$ $\frac{1}{4}(2+\sqrt{2\nu^{2}-2p\nu^{2}+2p^{2}\nu^{2}}).$
It can be shown that for all possible convex-mixing the maximum average success probability is the same as Eq.(62). Now, to be fully incompatible w.r.t convex-mixing $P(2,2,2)>\frac{3}{4},\,\,\text{which implies}\,\,\nu^{2}(p^{2}-p+1)>\frac{1}{2}$ (by Eq.(62)). Since the minimum value of $(p^{2}-p+1)$ is $\frac{3}{4},$ so for $\nu>\sqrt{\frac{2}{3}}$ the measurements are fully incompatible w.r.t convex-mixing.
In Observation 4, we have found that this kind of three qubit measurements with noise become fully incompatible w.r.t convex mixture for $\nu>\sqrt{\frac{2}{3}}$. So, we can conclude that when these three measurements are fully incompatible w.r.t convex mixing, quantum advantage is obtained in a suitable Random Access Code task.
[…]
Examples.— Consider the following three rank-one projective measurements acting on $\mathbb{C}^{3}$:
$X=\{|0_{x}\rangle\langle 0_{x}|,|1_{x}\rangle\langle 1_{x}|,|2_{x}\rangle\langle 2_{x}|\},$
$Y=\{|0_{y}\rangle\langle 0_{y}|,|1_{y}\rangle\langle 1_{y}|,|2_{y}\rangle\langle 2_{y}|\},$
and $Z=\{|0_{z}\rangle\langle 0_{z}|,|1_{z}\rangle\langle 1_{z}|,|2_{z}\rangle\langle 2_{z}|\},$ (63)
where,
$|0_{x}\rangle$ $=$ $\frac{1}{2}|0\rangle+\frac{1}{\sqrt{2}}|1\rangle+\frac{1}{2}|2\rangle,$
$|1_{x}\rangle$ $=$ $-\frac{1}{\sqrt{2}}|0\rangle+\frac{1}{\sqrt{2}}|2\rangle,$
$|2_{x}\rangle$ $=$ $\frac{1}{2}|0\rangle-\frac{1}{\sqrt{2}}|1\rangle+\frac{1}{2}|2\rangle,$
$|0_{y}\rangle$ $=$ $-\frac{\mathbbm{i}}{2}|0\rangle+\frac{1}{\sqrt{2}}|1\rangle+\frac{\mathbbm{i}}{2}|2\rangle,$
$|1_{y}\rangle$ $=$ $\frac{\mathbbm{i}}{\sqrt{2}}|0\rangle+\frac{\mathbbm{i}}{\sqrt{2}}|2\rangle,$
$|2_{y}\rangle$ $=$ $-\frac{\mathbbm{i}}{2}|0\rangle-\frac{1}{\sqrt{2}}|1\rangle+\frac{\mathbbm{i}}{2}|2\rangle,$
$|0_{z}\rangle$$\equiv$ $|0\rangle,|1_{z}\rangle\equiv|1\rangle,|2_{z}\rangle\equiv|2\rangle.$ (64)
We take the following convex mixture of the measurements $X$ and $Y$: $A=\{p|0_{x}\rangle\langle 0_{x}|+(1-p)|0_{y}\rangle\langle 0_{y}|,p|1_{x}\rangle\langle 1_{x}|+(1-p)|1_{y}\rangle\langle 1_{y}|,p|2_{x}\rangle\langle 2_{x}|+(1-p)|2_{y}\rangle\langle 2_{y}|\}$ with $0\leq p\leq 1$. We now consider the $(2,3,3)$ RAC game involving the two measurements $A$ and $Z$. The maximum average success probability $P(2,3,3)$ for this RAC task is given by Carmeli et al. (2020),
$P(2,3,3)=\frac{1}{18}\sum_{i,j=0}^{2}||A_{i}+Z_{j}||,$ (65)
which turns out to be greater than $\frac{2}{3}$ i.e. $P_{CB}(2,3,3)$ for all $p\in[0,1]$. Hence, the measurements $A$ and $Z$ are incompatible.
Next, consider the following measurement (taking an arbitrary convex mixture of $X$ and $Y$) $A_{i,j,k,l,m,n}=\{p|i_{x}\rangle\langle i_{x}|+(1-p)|j_{y}\rangle\langle j_{y}|,p|k_{x}\rangle\langle k_{x}|+(1-p)|l_{y}\rangle\langle l_{y}|,p|m_{x}\rangle\langle m_{x}|+(1-p)|n_{y}\rangle\langle n_{y}|\}$ with $0\leq p\leq 1$, $i,j,k,l,m,n\in\{0,1,2\}$, $i\neq k$, $k\neq m$, $m\neq i$ and $j\neq l$, $l\neq n$, $n\neq j$. Note that the aforementioned measurement $A$ is nothing but $A_{0,0,1,1,2,2}$ in the present notation. Following a similar calculation, it can be shown that any such $A_{i,j,k,l,m,n}$ is incompatible with $Z$ for all $p\in[0,1]$.
Similarly, taking an arbitrary convex combination of $X$ and $Z$ (or, $Y$ and $Z$), it can be shown that the new measurement is incompatible with $Y$ (or, $X$) for all $p\in[0,1]$.
A generic description for the semi-device independent witness of incompatibility.— Consider a measurement assemblage of $n$ measurements: $M=\{M_{0},M_{1},\ldots,M_{n-1}\}.$ We make $m$ partitions $S=\{S_{i}\}_{i=0}^{m-1}$, $\bigcup_{i}S_{i}=M,\,\text{and}\,S_{i}\bigcap S_{j}=\phi,\,\,\forall\,\,i,j\,\,\text{with}\,\,i\neq j.$ Our purpose is to operationally witness the incompatibility of $m$ measurements produced from the convex mixing of $n$ measurements from each of the $m$ partitions. We can construct a $(m,d,d)$ RAC game where Alice has an $m$ dit input message, viz., $x=(x_{0},x_{1},\ldots,x_{m-1})$. Depending upon the message, she encodes it in a qudit and sends it to Bob, who on the other hand, gets input $y\in\{0,1,\ldots,m-1\}$ and accordingly, he has to predict the value of the corresponding bit $x_{y}$. He performs the measurement which is obtained by the convex mixtures of the measurements from the partition $S_{y}$ and declares the outcome of the measurement. Now, if the success probability $P(m,d,d)$ is greater than $P_{CB}(m,d,d)$ Saha et al. (2023) for all possible convex mixtures, we can operationally detect that the $m$ partitioned measurements are incompatible under convex mixing.
V Conclusions
Measurement incompatibility is the spectacle in quantum theory that a set of measurements cannot be performed jointly on arbitrary systems. It is one of the fundamental ingredients for non-classical correlations and merits of quantum information science. Incompatibility offers a complex structure with different layers as the number of measurements and the dimension on which the measurements act increases. Thus, understanding the different levels of incompatibility with respect to elementary classical operations is of paramount importance both from the foundational and practical perspectives. In this work, we have considered the two most general classical operations viz., coarse-graining of different outcomes of measurements, and convex-mixing of different measurements.
Through our present analysis we have investigated the different levels of incompatibility arising under the above classical operations. Since environmental effects are ubiquitous in practical scenarios, here we have investigated the tolerance thresholds for maintaining measurement incompatibility against noise under classical operations. Furthermore, we have developed a method to operationally witness different levels of measurement incompatibility in the device-independent framework involving Bell-type experiments, and also in the semi-device-independent framework involving prepare-and-measure experiments.
Several examples have been provided to illustrate the efficacy of the operational witnesses proposed here. Our results are particularly useful for the purpose of comparing measurements in terms of their degree of incompatibility. For example, measurements that remain fully incompatible with respect to coarse-graining of outcomes (or convex-mixing of measurements) show stronger incompatibility compared to those that are not fully incompatible, thus facilitating legitimate choice of measurements in an information-processing task where a high degree of incompatibility is required.
Our present study motivates future work on several open directions. The condition for full incompatibility for two projective measurements with respect to coarse-graining of outcomes can be extended for more number of projective measurements. Similarly, the criterion for full incompatibility of projective measurements with respect to convex mixing could be generalized for any set of measurements. Further investigation would be required to see if device-independent witnesses could be formulated for more than three outcome measurements with respect to coarse-graining and more than three measurements with repect to convex mixing. It will be interesting to look for examples where the full incompatibility with respect to coarse-graining can be inferred in a device-independent way from a single experimental statistics.
VI Acknowledgements
[…]

### v2

@@
 Abstract
-Measurement incompatibility has proved to be an important resource for information processing tasks.
-In this work, we analyze various levels of incompatibility of measurement sets.
-We provide operational classification of measurement incompatibility with respect to two elementary classical operations, viz., coarse-graining of measurement outcomes and convex mixing of different measurements.
-We derive analytical criteria for determining when a set of projective measurements is fully incompatible with respect to coarse-graining or convex mixing.
-Robustness against white noise is investigated for mutually unbiased bases that can sustain full incompatibility.
-Furthermore, we propose operational witnesses for different levels of incompatibility subject to classical operations, using the input-output statistics of Bell-type experiments as well as experiments in the prepare-and-measure scenario.
+Measurement incompatibility has proved to be an important resource for information processing.
+In this work, we present an operational approach that leverages classical operations on the inputs and outputs of measurement devices to explore different layers of incompatibility among the measurements performed by the device.
+We study classifications of measurement incompatibility with respect to two types of classical operations, viz., coarse-graining of measurement outcomes and disjoint-convex-mixing of different measurements.
+We derive analytical criteria for determining when a set of projective measurements is fully incompatible with respect to coarse-graining or disjoint-convex-mixing.
+Robustness against white noise for different layers of incompatibility for mutually unbiased bases is investigated.
+Furthermore, we study operational witnesses for incompatibility subject to these classical operations, using the input-output statistics of Bell-type experiments as well as for experiments in the prepare-and-measure scenario.
 I introduction
-Measurement incompatibility is a concept relating observables which cannot be measured jointly with arbitrary accuracy Gühne et al.
+Measurement incompatibility is a concept related to observables that cannot be measured jointly with arbitrary accuracy Gühne et al.
 (2023).
-It is a purely quantum effect of which the most well-known example concerns the position and momentum of a quantum particle.
+It is purely a quantum effect, of which the most well-known example concerns the position and momentum of a quantum particle.
 Being a fundamental concept of quantum theory, it takes a pivotal role in explaining several quantum phenomena, such as, Bell-nonlocality Fine (1982a); Wolf et al.
-(2009), Einstein–Podolsky–Rosen steering Uola et al.
+(2009a), Einstein-Podolsky-Rosen steering Uola et al.
 (2014); Quintino et al.
@@
 (2018).
-Measurement incompatibility finds applications in various quantum information processing tasks.
-Recently, it has been shown that measurement incompatibility is necessary for quantum advantage in any one-way communication task Saha et al.
+Apart from the foundational significance of incompatible measurements, they have been proven to be an important resource for various information processing tasks Skrzypczyk et al.
+(2019); Buscemi et al.
+(2020); Carmeli et al.
 (2019).
-The significance of measurement incompatibility in various quantum effects calls for its in-depth characterization.
+Recently, It is also known that measurement incompatibility is necessary for quantum advantage in any one-way communication task Saha et al.
+(2023).
+The significance of measurement incompatibility in various operational tasks calls for its in-depth characterization.
 Towards this direction, a classification of measurement incompatibility with respect to projection onto subspaces has been recently performed Uola et al.
+(2021), also different hierarchy of incompatible measurements has been given based on the number of copies of quantum states required Carmeli et al.
+(2016), and the simulability of their statistics produced in the experiments Filippov et al.
+(2018); Egelhaaf et al.
 (2024).
-In the present work, our aim is to classify measurement incompatibility with respect to certain basic classical operations.
-From an operational perspective, here we employ two such operations, namely coarse-graining of measurement outcomes and convex mixing of measurement settings.
-Coarse-graining of measurement outcomes arises naturally Toscano et al.
-(2018) in several instances, such as, for example, quite obviously in measurements on continuous variable system.
-Though the eigen spectra of the observables are infinite-dimensional taking continuous values, real-world experimental devices are limited by finite precision, leading to the measurement outcomes taking a finite number of discrete values.
+In the continuous variable system, the verification of incompatibility using phase-space quasiprobability distributions has been studied in Ref.Rahimi-Keshari et al.
+(2021).
+In the present work, our objective is to classify measurement incompatibility in an operational approach that does not involve the details of a theory.
+To put it differently, we address how various degrees of incompatibility can be assessed solely by executing basic classical operations on the inputs or outputs of these measurements.
+From an operational perspective, when the concerned measurement devices are black boxes with no control over the internal workings, one can still realize different measurements by performing suitable classical operations to manipulate the statistics of the measurements.
+Here, we employ two such operations: coarse-graining of measurement outcomes (classical post-processing performed on the outcomes of a measurement) and convex-mixing of measurement settings (classical operation performed on the inputs).
+One can also consider classical operations performed on both inputs and outputs.
+Similar types of classical operations were also introduced in Haapasalo et al.
+(2012), nevertheless, the study of incompatibility under these classical operations was lacking.
+The motivation of this present work is to fill this crucial gap in the literature.
+Coarse-graining of measurement outcomes arises naturally in several instances, for example, in measurements on continuous variable systems Toscano et al.
+(2018).
+Though the eigen spectra of the observables are infinite-dimensional and continuous, real-world experimental devices are limited by finite precision, leading to the measurement outcomes taking a finite number of discrete values.
 This inaccuracy in the recording of measurement outcomes is manifested in the coarse-graining of measurement outcomes, which is inevitable in practice.
-Coarse-graining has also been employed to study the phenomenon of quantum to classical transition.
-It is observed that quantum phenomena may disappear due to imprecision of measurement outcomes Kofler and Brukner (2007); Rudnicki et al.
+On the other hand, device imperfection may also lead to the measurement device performing a set of measurements probabilistically, instead of always performing the desired particular measurement.
+In such a case, a convex-mixing of the given set of measurements arises effectively D’Ariano et al.
+(2005).
+By leveraging these classical operations on the measurement device, we can learn the finer details about the measurements i.e.
+the different layers of incompatibility, which can give a measure of the degree of incompatibility of a set of measurements.
+This can set a benchmark for legitimately choosing incompatible measurements for an information processing task.
+Our study is motivated to address how one may compare the degree of incompatibility between two different sets of measurements subjected to the aforementioned classical operations.
+For instance, if the first set remains incompatible for every possible non-trivial coarse-graining of the measurement outcomes, but the second set becomes compatible for a certain coarse-graining, it follows that the first set of measurements exhibits stronger incompatibility compared to the second one.
+A similar argument holds for the case of convex-mixing of measurements also.
+In this work, we establish analytical criteria for determining when a pair of projective measurements are fully incompatible, i.e., remain incompatible under all possible coarse-grainings of measurement outcomes.
+We further analyze the full incompatibility of a set of three qubit measurements under all possible convex-mixing as well.
+Within the context of our present study, noise is reflected in degrading the incompatibility properties of various measurement sets Designolle et al.
+(2019a).
+We compute the critical noise threshold below which the mutually unbiased bases measurements remain incompatible under the above-mentioned classical operations.
+As incompatible measurements are useful for various information processing tasks, any device claiming to produce incompatible measurements must be certified before using it in an experiment.
+Verification of incompatibility of the measurements is possible from the input-output measurement statistics obtained from the device without knowing its internal functioning.
+This can be done in two ways one is a device-independent way utilizing Bell-type experiments Wolf et al.
+(2009b); Bene and Vértesi (2018); Hirsch et al.
+(2018) and another is a semi-device-independent way inspired by the standard prepare-and-measure scenario de Gois et al.
+(2021); Saha et al.
+(2023); Carmeli et al.
+(2020).
+In the present work, we investigate the issue of certification of different layers of measurement incompatibility under the introduced classical operations both from the device-independent and the semi-device-independent perspective.
+As both Bell-type experiments and prepare-and-measure experiments Zhang et al.
+(2024); Tavakoli et al.
+(2015) are viable with the current technology, by performing our introduced classical operations on those experiments we can operationally certify the different layers of incompatibility thus allowing us to know the subtleties of incompatibility about the measurement device which is a new and interesting offshoot of our work.
+Our paper is structured as follows.
+In Sec.(II), the significance of these classical operations in an operational paradigm is discussed along with the physical ground for these classical operations.
+In Sec.
+(III), we define a hierarchy in the incompatibility of measurements under these classical operations.
+In Sec.
+(IV), we study the noise robustness for various levels of incompatibility of measurements subjected to the above classical operations.
+In Sec.
+(V), we define operational witnesses of incompatibility of measurements under these classical operations and furnish examples to study their performance in device-independent and semi-device-independent frameworks.
+Concluding remarks are presented in Sec.(VI).
+II Operational approach to classifying incompatibility
+Consider a measurement device where, when a physical system is probed, we can choose measurement settings via a classical parameter, denoted by $x$.
+This parameter serves as the input to the device, and the device produces the measurement outcome $z$, corresponding to that input.
+Suppose there are $n$ possible inputs, i.e., $x\in[n]$, and for each input, there are $d$ possible outcomes, i.e., $z\in[d]$, where $[k]$ represents the set $\{0,\cdots,k-1\}$ for any natural number $k$.
+For simplicity, we assume that the number of outcomes for each measurement is the same.
+Without loss of generality, this assumption holds, as any measurement with fewer than $d$ outcomes can be treated as having $d$ outcomes by considering some outcomes as never occurring.
+By probing different system preparations, we can gather input-output statistics from the device.
+In quantum mechanics, measurements are described by Positive Operator Valued Measures (POVMs), which consist of positive semi-definite operators that sum up to the identity operator.
+Let us represent the measurements realized in the device by the set of operators $\{M_{z|x}\}_{z,x}$, where $x$ indexes different measurements, and $z$ denotes the corresponding outcomes.
+A set of measurements is said to be compatible if there exists a parent POVM, $G_{\lambda}$, and classical post-processing $\{p(z|x,\lambda)\}$ for each $x$ such that
+$\forall z,x,\quad M_{z|x}=\sum_{\lambda}p(z|x,\lambda)G_{\lambda},$ (1)
+where $0\leq p(z|x,\lambda)\leq 1,$ and $\sum_{z}p(z|x,\lambda)=1,$ for all $x,\lambda$ Gühne et al.
+(2023).
+As a special case, if the operators are projectors, then the two measurements are jointly measurable when their corresponding operators commute.
+In an operational paradigm, we do not have direct control over the internal workings of the measurement device.
+However, we can manipulate the classical inputs and outputs to realize different measurements.
+Figure 1: Classical operations on the outputs (left) and classical operations on the inputs (right) to realize a new set of measurements are depicted.
+II.1 Classical operations on outputs
+In general, one can modify the outcomes of a measurement by applying a post-processing operation or stochastic map (see FIG.
+1).
+For a given input $x$, this operation is mathematically described by a probability distribution $\mathcal{C}_{x}:=\{c_{x}(\tilde{z}_{x}|z)\}$, $c_{x}(\tilde{z}_{x}|z)\geqslant 0$ and $\sum_{\tilde{z}_{x}}c_{x}(\tilde{z}_{x}|z)=1$.
+Here $\tilde{z}_{x}\in[\tilde{d}_{x}]$ denotes the outcome of the new measurement after the operation, and $\mathcal{C}_{x}$ is the stochastic map with $c_{x}(\tilde{z}_{x}|z)$ is the probability of producing outcome $\tilde{z}_{x}$ when outcome $z$ is obtained for the measurement $x$.
+In general, one can apply different operations for different $x$.
+By applying different operations for each $x$, one can generate a new set of measurements described by the operators $\{\tilde{M}_{\tilde{z}_{x}|x}\}_{\tilde{z}_{x},x}$ as follows:
+$\tilde{M}_{\tilde{z}_{x}|x}=\sum_{z}c_{x}(\tilde{z}_{x}|z)M_{z|x}\,\,,$ (2)
+for each $x.$ If the initial set of measurements is incompatible, we can ask whether this new set is incompatible or not.
+The set of all possible operations $\{\mathcal{C}_{x}\}$ is infinite.
+Furthermore, certain operations, such as generating random outputs, will always result in a set of measurements that is compatible, regardless of the original measurements.
+To avoid such intricacies, we focus on a particular class of operations where $c_{x}(\tilde{z}_{x}|z)\in\{0,1\}$ and $\tilde{d}_{x}<d$.
+These classical operations involve coarse-graining or relabelling/permuting outcomes (or a combination of both).
+This class contains a finite number of operations, and interestingly, any classical operation on outputs can be expressed as convex combinations of this class of operations.
+More importantly, certain quantum measurements remain incompatible even after all possible operations of this kind.
+We will refer to this class of operations as coarse-graining.
+Coarse-graining typically refers to treating more than one outcome as equivalent, i.e., multiple outcomes are clubbed together to form a single outcome.
+As a result, the effective number of outcomes is reduced.
+II.2 Classical operation on inputs
+We can also apply classical operations to the inputs to generate a new set of measurements, each of which is a convex combination of the initial measurements (see FIG.
+1).
+This classical operation is represented by a probability distribution $\{e(x|\tilde{x})\}_{x,\tilde{x}}$, where $e(x|\tilde{x})\geqslant 0$ and $\sum_{x}e(x|\tilde{x})=1$ for each $\tilde{x}\in[m]$.
+Here, $e(x|\tilde{x})$ refers to the convex weightage of appearing measurement $x$ in the new measurement labelled by $\tilde{x}$.
+The new set of $m$ measurements, defined by the operators $\{\tilde{M}_{z|\tilde{x}}\}_{z,\tilde{x}}$, is given by:
+$\tilde{M}_{z|\tilde{x}}=\sum_{x}e(x|\tilde{x})M_{z|x},$ (3)
+for every $z$.
+If the same convex mixing is applied uniformly to all inputs, i.e., $e(x|\tilde{x})$ is the same for all $\tilde{x}$, the resulting measurements will be identical and hence compatible.
+To avoid such cases, we introduce a particular class of convex mixing, which we will refer to as disjoint-convex-mixing of measurements.
+In disjoint-convex-mixing, no initial measurement appears in the convex combination of more than one new measurement.
+Mathematically, for every $z$, only one of the numbers from the set $\{e(x|\tilde{x})\}_{\tilde{x}}$ is non-zero, and others are zero, which can be expressed as:
+$\forall x,\ \max_{\tilde{x}}\{e(x|\tilde{x})\}=\sum_{\tilde{x}}e(x|\tilde{x}).$ (4)
+In this work, we focus on disjoint-convex-mixing for classifying measurement incompatibility.
+We will point out later that there are measurements that remain incompatible after all possible disjoint-convex-mixing.
+II.3 Classical operations on both inputs and outputs
+It is also possible to construct new measurements by combining operations on both inputs and outputs.
+This can be done in two ways - first applying an operation on the inputs followed by the outputs, or vice versa.
+In the former case, the operators $\{\tilde{M}_{\tilde{z}_{\tilde{x}}|\tilde{x}}\}_{\tilde{z}_{\tilde{x}},\tilde{x}}$ defining the new measurements are
+$\tilde{M}_{\tilde{z}_{\tilde{x}}|\tilde{x}}=\sum_{z}\sum_{x}c(\tilde{z}_{x}|z)e(x|\tilde{x})M_{z|x}\ ,$ (5)
+where $\{e(x|\tilde{x})\}$ and $\{c(\tilde{z}_{x}|z)\}_{x}$ denote the classical operation on inputs and outputs, respectively.
+In the latter case, the set of operators $\{\tilde{M}_{\tilde{z}|\tilde{x}}\}$ are
+$\tilde{M}_{\tilde{z}|\tilde{x}}=\sum_{x}\sum_{z}e(x|\tilde{x})c(\tilde{z}|z)M_{z|x}\ ,$ (6)
+where $\{e(x|\tilde{x})\}$ and $\{c(\tilde{z}|z)\}$ denote the classical operation on inputs and outputs, respectively.
+Here again, without loss of generality, we consider the number of outcomes after the operations on the outcomes to be the same, which is denoted by $\tilde{z}$.
+It is important to note that these two cases are not equivalent, as the stochastic operations on the outputs apply to different sets of measurements in each case.
+In summary, one can explore the effects of various classical operations on both inputs and outputs to determine whether the resulting measurements remain incompatible.
+However, this work specifically focuses on two key classes of operations: coarse-graining of outputs and disjoint-convex-mixing of inputs.
+In addition to the operational relevance of the two types of operations mentioned above, their significance for practical implementation is discussed in the following subsection.
+II.4 Physical motivations for considering the two specific classical operations
+Let us now elaborate further the motivations for considering the two aforementioned specific classical operations on a set of measurements.
+Coarse-graining of measurement outcomes is a natural consequence of an observer’s limitation in a practical scenario involving multi-outcome measurements.
+For example, consider the measurement of the spin-$z$ component (associated with the operator $J_{z}$) of the spin-$j$ system, where $j$ is very large.
+The possible outcomes of this measurement are nothing but the eigenvalues of $J_{z}$, denoted by $m\in\{-j,-j+1,\cdots,j\}$.
+In practice, this type of measurement is performed using a Stern-Gerlach type experiment.
+In such a practical situation, the concept of “neighbouring outcomes” arises Kofler and Brukner (2007).
+For example, the outcomes $m$ and $m+1$ are neighbouring outcomes in the real configuration space in a Stern-Gerlach type experiment.
+For large $j$, it is almost impossible for a device with finite precision to resolve these neighbouring outcomes in the observation screen – giving rise to coarse-graining of measurement outcomes.
+This type of practical limitation is not only limited to large spins but also applicable to other multi-outcome measurements and measurements of continuous variables (e.g., position, momentum Das et al.
+(2024)).
+Note that coarse-graining of measurement outcomes has been invoked in the context of explaining classical limits of quantum mechanics Kofler and Brukner (2007); Rudnicki et al.
 (2012); Jeong et al.
-(2014); Mal and Majumdar (2016).
-Device imperfection may also lead to the measurement device performing probabilistically a set of measurements, instead of always performing the desired particular measurement.
-In such a case, a convex mixing of the given set of measurements arises effectively.
-As incompatibility of measurements is a quantum concept, it is interesting to examine how this property behaves under such elementary classical operations.
-Our study is motivated by the question as to how one may compare the degree of incompatibility between two different sets of measurements.
-For instance, if the first set remains incompatible for every possible non-trivial coarse-graining of the measurement outcomes, but the second set becomes compatible for a certain coarse-graining, it follows that the first set of measurements exhibits stronger incompatibility compared to the second one.
-A similar argument holds for the case of convex-mixing of measurements too.
-As incompatible measurements are useful for performing various information processing tasks, given a measurement device claiming to produce incompatible measurements, we must certify it before using it in an information processing task.
-Verification of incompatiblity of the measurements is possible from the input-output measurement statistics obtained from the device without knowing its internal functioning.
-Device-independent protocols are conceptually most powerful, relying only on the input-output statistics Brunner et al.
-(2014), with a wide range of applications from quantum cryptography Ekert (1991); Acín et al.
-(2007) to communication complexity Buhrman et al.
-(2010), self-testing Mayers and Yao (2004); Šupić and Bowles (2020); Goswami et al.
-(2018); Bian et al.
-(2020) and incompatibility witnesses Quintino et al.
+(2014); Mal and Majumdar (2016); Mal et al.
+(2016); Mukherjee et al.
 (2019).
-However, they require shared entanglement, an expensive resource, and prohibition of communication between the involved parties.
-On the other hand, semi-device-independent protocols are inspired by the standard prepare-and-measure scenario, with an additional constraint on the dimension of the communicated quantum states Van Himbeeck et al.
-(2017); Moreno et al.
-(2021).
-They have been applied in quantum key distribution Pawłowski and Brunner (2011), quantum networks Bowles et al.
-(2015), randomness certification Passaro et al.
-(2015), random access codes Li et al.
-(2012), and certification of meassurements Maity et al.
-(2021); Das et al.
-(2022).
-In the present work we investigate the issue of certification of measurement incompatibility subjected to the classical operations of coarse-graining and convex mixing of measurements from both the device-independent and the semi-device-independent perspective.
-An essential ingredient in the performance of quantum information protocols is the presence of noise.
-Within the context of our present study, noise is reflected in the degrading of incompatibility properties of various measurement sets.
-We examine the critical amount of robustness of the noisy measurement above which the measurements remain incompatible under the mentioned classical operations.
-We then study incompatibility witnesses in the device-independent and the semi-device-independent scenarios.
-Our paper is structured as follows.
-In Sec.
-(II) we present the basic mathematical ingredients for classifying measurement incompatibility under the two operations of coarse-graining of outcomes and convex-mixing of measurements, respectively.
-In Sec.
-(III) we study the robustness under noise for various levels of incompatibility subjected to the above classical operations.
-In Sec.
-(IV) we define operational witnesses of incompatibility, and furnish examples to study their performance in both device-independent and the semi-device-independent frameworks.
-Concluding remarks are presented in Sec.
-(V).
-II Classifying measurement incompatibility under elementary classical operations
-In the case of projective measurements, the observables whose corresponding operators commute are jointly measurable.
-However, this is not true in the case of general quantum measurements.
-The most general quantum measurement is described by Positive Operator Valued Measures (POVM), which is a set of operators, $\{O_{\alpha}\}$ with $0\leq O_{\alpha}\leq\mathbbm{1}$, and $\sum_{\alpha}O_{\alpha}=\mathbbm{1}$.
-A set of measurements defined by $\{M_{z_{x}|x}\}$ (here $x$ corresponds to the measurement, $z_{x}$ denotes the outcomes of the measurement labeled by $x$) is compatible if there exists a parent POVM, $G_{\lambda}$ and classical post-processing for each $x$ given by $\{p(z_{x}|x,\lambda)\}$ such that
-$\forall z_{x},x,\quad M_{z_{x}|x}=\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda},$ (1)
-where $0\leq p(z_{x}|x,\lambda)\leq 1,\sum_{z_{x}}p(z_{x}|x,\lambda)=1\forall x,\lambda$    Gühne et al. (2023).
-II.1 Coarse-graining of outcomes
-Coarse-graining of the outcomes of a measurement is a concept where some of the outcomes are treated on the same footing, i.e., different outcomes are clubbed together to form a single outcome.
-After the coarse-graining of the outcomes, the effective number of outcomes reduces.
-Coarse-graining may occur in practice due to measurement errors.
-Additionally, coarse-graining is used in when one does not require a finer description of the values of the measurement outcomes, but rather a broad description suffices to understand the nature of some physical property.
-Let us introduce the notation $[k]:=\{0,\cdots,k-1\}$ for any natural number $k.$ Consider a $\overline{d}$ outcome measurement $\{M_{\overline{z}}\}$ and an arbitrary coarse-graining that yields a $d$ outcome measurement $\{M_{z}\}$, where $\overline{z}\in[\overline{d}]\,\,\text{and}\,\,z\in[d]$.
-The coarse-graining can be defined by conditional probabilities $\{c(z|\overline{z})\}$ such that
-$M_{z}=\sum_{\overline{z}}c(z|\overline{z})M_{\overline{z}},$ (2)
-and $c(z|\overline{z})\in\{0,1\}$.
-Trivial coarse-graining is the one where the measurement becomes single outcome, that is, $c(z|\overline{z})=1,\forall\overline{z}$.
-Two incompatible measurements may not necessarily remain incompatible after certain coarse-graining of their outcomes.
-Definition 1 (Fully incompatible measurements w.r.t coarse-graining).
-A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is fully incompatible with respect to (w.r.t) coarse-graining if they remain incompatible after all possible nontrivial coarse-graining.
-That is, if the resultant set of measurements given by
-$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},$ (3)
-after all possible sets of nontrivial coarse-graining $\{c_{x}(z_{x}|\overline{z}_{x})\}$ pertaining to every setting $x$ is incompatible, then we call them fully incompatible.
-Definition 2 ($d$-incompatible measurements w.r.t coarse-graining).
-A set of measurements $\{M_{\overline{z}_{x}|x}\}$ is $d$-incompatible w.r.t coarse-graining if they remain incompatible after all possible nontrivial coarse-graining that give rise to at least $d$ outcome measurements.
-In other words, if the resultant set of measurements
-$M_{z_{x}|x}=\sum_{\overline{z}_{x}}c_{x}(z_{x}|\overline{z}_{x})M_{\overline{z}_{x}|x},\ z_{x}\in[k_{x}],k_{x}\geqslant d,\forall x$ (4)
-after all possible coarse-graining resulting in at least $d$ outcome measurements is incompatible, then we call them $d$-incompatible.
-For example, the four outcome rank-one projective measurement pair, defined by the following (un-normalized) vectors
-$M=\Big{\{}|0\rangle,|1\rangle,|2\rangle,|3\rangle\Big{\}}$ (5)
+On the other hand, convex mixing of measurements is also inevitable in a practical situation.
+To explain it in more detail, let us consider the example of measuring $\sigma_{z}$ (Pauli spin observable) on a qubit.
+In practice, the observable to be measured ($\sigma_{z}$ in the present example), defined by the relative direction of the inhomogeneous magnetic field in the Stern-Gerlach apparatus with respect to the direction of the incoming beam of spin-$1/2$ particles, may not be kept fixed in all experimental runs.
+Consequently, instead of $\sigma_{z}$, $\vec{\sigma}.\hat{n}$ will be measured ($\vec{\sigma}.\hat{n}$ = $\sigma_{x}n_{x}+\sigma_{y}n_{y}+\sigma_{z}n_{z}$ with $\sigma_{x}$, $\sigma_{y}$, $\sigma_{z}$ are three Pauli operators and $n_{x}^{2}+n_{y}^{2}+n_{z}^{2}=1$), where $\hat{n}$ will be different in different runs (all such $\hat{n}$ should be close to $\hat{z}$).
+Effectively, this will give rise to convex mixing of different measurements.
+Hence, these operations are extremely relevant in practical scenarios.
+The various layers of incompatibility and their operational implications under these operations are discussed in the subsequent sections.
+III Classifying measurement incompatibility under coarse-graining and disjoint-convex-mixing
+III.1 Coarse-graining of outcomes
+Recall that the coarse-graining of a set of $d$-outcome measurements $\{M_{z|x}\}$ produces new measurements defined by
+$\tilde{M}_{\tilde{z}_{x}|x}=\sum_{z}c_{x}(\tilde{z}_{x}|z)M_{z|x},\ \text{ where }\ c_{x}(\tilde{z}_{x}|z)\in\{0,1\}.$ (7)
+Before moving forward, let us define what we mean by trivial coarse-graining.
+If the coarse-graining results in a measurement where one of the outcomes, say $\tilde{z}_{x}$, always occurs, that is, $\tilde{M}_{\tilde{z}_{x}|x}=\mathbbm{1}$, we refer that operation as trivial coarse-graining for the measurement labelled by $x$.
+The fact that two incompatible measurements may not necessarily remain incompatible after certain coarse-graining motivates the following definition of full incompatibility with respect to (w.r.t.) coarse-graining.
+Definition 1 (Fully incompatible measurements w.r.t. coarse-graining).
+A set of measurements $\{M_{z|x}\}$ is fully incompatible w.r.t.
+coarse-graining if they remain incompatible after all possible nontrivial coarse-graining.
+That is if the resultant set of measurements given by Eq.
+(7) after all possible sets of nontrivial coarse-graining is incompatible, then we call them fully incompatible.
+Note that, the coarse-graining can be different for different settings $x$.
+Definition 2 ($k$-incompatible measurements w.r.t. coarse-graining).
+A set of measurements $\{M_{z|x}\}$ is $k$-incompatible w.r.t.
+coarse-graining if they remain incompatible after all possible nontrivial coarse-graining that gives rise to at least $k$ outcome measurements.
+In other words, if the resultant set of measurements given by Eq.
+(7), where $\tilde{z}_{x}\in[d_{x}]$ and $d_{x}\geqslant k$ for all $x,$ is incompatible, then we call them $k$-incompatible.
+For example, the three outcome rank-one projective measurement pair, defined by the following (un-normalized) vectors
+$M=\big{\{}|0\rangle,|1\rangle,|2\rangle\big{\}}$ (8)
 and
-$N=\Big{\{}|0+1\rangle,|0-1\rangle,|2+3\rangle,|2-3\rangle\Big{\}}$ (6)
-is 3-incompatible, where the notation $|0+1\rangle=|0\rangle+|1\rangle,$ and similarly for others.
+$N=\big{\{}|0\rangle+|1\rangle,|0\rangle-|1\rangle,|2\rangle\big{\}}$ (9)
+is 3-incompatible, but not 2-incompatible since coarse-graining of the first two outcomes of these measurements yields compatible measurements.
 Observation 1.
-A set of fully incompatible measurements is equivalent to 2-incompatible measurements w.r.t coarse-graining.
-If a set of measurements is 2-incompatible, then it implies that the set remains incompatible after all possible coarse-graining of the outcomes such that the number of outcomes of each measurement in the newly formed set of measurements is greater than or equal to two.
-Also, the lowest number of outcomes of measurement is two for a nontrivial coarse-graining.
-Furthermore, if a set of measurements is $d$-incompatible, then, by definition, it is $n$-incompatible as well, where $n>d$, but the reverse is not true.
-This proves Observation 1.
+A set of fully incompatible measurements is equivalent to 2-incompatible measurements w.r.t. coarse-graining.
+Proof.
+[…]
 Observation 2.
-[…]
+Consider two projective measurements, defined by $\{P_{i}\},\,\,\text{and}\,\,\{Q_{j}\}$, where $i\in[d]$ and $j\in[d^{\prime}]$.
+Let $\{\mathcal{M}_{k}\}_{k}$ be the set of all proper subsets of $[d]$, and $\{\mathcal{N}_{l}\}_{l}$ be the set of all proper subsets of $[d^{\prime}]$.
+Then these two measurements are fully incompatible w.r.t.
+coarse-graining if and only if
 $\left[\sum_{i\in\mathcal{M}_{k}}P_{i},\sum_{j\in\mathcal{N}_{l}}Q_{j}\right]\neq 0,\ \forall k,l.$ (10)
@@
 Theorem 1.
-The following condition is necessary but not sufficient for two rank-one projective measurements, defined by $\{|\psi_{i}\rangle\}$ and $\{|\phi_{j}\rangle\}$, to be fully incompatible w.r.t coarse-graining:
+The following condition is necessary but not sufficient for two rank-one projective measurements of dimension $\geq 4$, defined by $\{|\psi_{i}\rangle\}$ and $\{|\phi_{j}\rangle\}$, to be fully incompatible w.r.t.
+coarse-graining:
 $\langle\psi_{i}|\phi_{j}\rangle\neq 0,\ \forall i,j.$ (11)
-The above condition is necessary and sufficient for two 3-dimensional rank-one projective measurements.
+However, the above condition is necessary and sufficient for two 3-dimensional rank-one projective measurements.
 Proof.
-First, note that for sharp measurement compatibility and commutativity are equivalent Heinosaari and Wolf (2010).
+First, note that for sharp measurement, compatibility and commutativity are equivalent Heinosaari and Wolf (2010).
 Suppose $\exists i,j,$ such that $\langle\psi_{i}|\phi_{j}\rangle=0$.
@@
 […]
-II.2 Convex mixing of measurements
-The concept of convex mixing of measurements may be best understood by considering an example: suppose we have a measurement device that produces three different measurements $(M_{0},M_{1}\,\,\text{and}\,\,M_{2})$ and has two input switches.
-If the first switch is ON it performs measurement $M_{0}$ with probability $q$ and $M_{1}$ with probability $(1-q)$ and for the second switch, it always performs $M_{2}.$ Thus, in principle, the device produces two measurements: $M_{2}$ and a convex mixture of $M_{0}$ and $M_{1}.$
-Suppose $M_{0}=\{M_{z|0}\},M_{1}=\{M_{z|1}\},M_{2}=\{M_{z|2}\}$ such that they have same number of outcomes $z\in[d]$.
-Now the measurement that is realized by a convex combination of $M_{0}$ and $M_{1}$ with some weightage $q\,\,\text{and}\,\,1-q$, where $q\in[0,1]$is defined as
-$M_{(0,1,\pi)}=\{qM_{z|0}+(1-q)M_{\pi(z)|1}\},$ (11)
-where $\pi(z)$ is any permutation on outcome $z$.
-Here $\pi(z)$ is a bijective function from the set $z\in[d]$ to itself.
-Even if $M_{2}$ is incompatible with $M_{0}$ and $M_{1}$ separately, $M_{2}$ is not necessarily incompatible with $M_{(0,1,\pi)}$ for all values of $q$.
+III.2 Disjoint-convex-mixing of measurements
+Consider a device that implements three different measurements, $M=\{M_{z}\}_{z},N=\{N_{z}\}_{z}$, and $R=\{R_{z}\}_{z}$, all having the same number of outcomes $z\in[d]$.
+For the first input, it performs measurement $M$ with probability $q$ and $N$ with probability $(1-q)$.
+For the second input, it always performs measurement $R.$ Thus, the new measurement $Q_{(M,N)}$, realized through the convex-mixing, is
+$Q_{(M,N)}=\{qM_{z}+(1-q)N_{z}\}_{z},$ (14)
+where $q\in[0,1]$ is the weightage of the disjoint-convex-mixing.
+Even if $R$ is incompatible with $M$ and $N$ separately, $R$ is not necessarily incompatible with $Q_{(M,N)}$ for all values of $q$.
+In a similar way, we introduce the notion of full incompatibility w.r.t.
+disjoint-convex-mixing.
 Definition 3.
-[…]
-Consider three unbiased qubit measurements,
-$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\vec{n_{i}}\cdot\vec{\sigma}\right),\quad i=0,1,2,$ (12)
-and $||\vec{n_{i}}||\leqslant 1$. These three measurements are fully incompatible w.r.t convex mixtures if and only if (see eq.(7) of Gühne et al. (2023)),
-$||\vec{n}_{i}+q\vec{n}_{j}\pm(1-q)\vec{n}_{k}||+||\vec{n}_{i}-q\vec{n}_{j}\mp(1-q)\vec{n}_{k}||>2,$ (13)
+Three measurements $M,N\,\,\text{and}\,\,R$ are fully incompatible w.r.t.
+disjoint-convex-mixing if each of the pairs, $M$ and $Q_{(N,R)},$ $N$ and $Q_{(M,R)},$ $R$ and $Q_{(M,N)},$ are incompatible for all values of $q$, where $Q_{(\ ,\ )}$ is defined in (14).
+Consider three unbiased qubit binary outcome measurements $\{M_{0},M_{1}\},\{N_{0},N_{1}\},\{R_{0},R_{1}\}$,
+$M_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\,\,\vec{n_{0}}\cdot\vec{\sigma}\right),$
+$N_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\,\,\vec{n_{1}}\cdot\vec{\sigma}\right),$
+$R_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\,\,\vec{n_{2}}\cdot\vec{\sigma}\right),$ (15)
+with $z=0,1$ and $||\vec{n_{i}}||\leqslant 1\,\,\text{where}\,\,i\in\{0,1,2\}$.
+A necessary and sufficient criterion for the incompatibility of two unbiased binary-outcome qubit measurements is given by Busch in Busch (1986) (also, see eq.(7) of Gühne et al.
+(2023)).
+By applying this criterion, we find that the above three measurements (III.2) are fully incompatible w.r.t.
+disjoint-convex-mixing if and only if,
+$||\vec{n}_{i}+q\vec{n}_{j}+(1-q)\vec{n}_{k}||+||\vec{n}_{i}-q\vec{n}_{j}-(1-q)\vec{n}_{k}||>2,$ (16)
 for all $q$, and for all $(i,j,k)\in\{(0,1,2),(1,2,0),(2,0,1)\}$.
-Observation 3.
-If three qubit measurements (12) are such that $\vec{n}_{i}$ are in the same plane of the Bloch sphere, then they are not fully incompatible w.r.t convex mixtures.
+Theorem 2.
+If three-qubit measurements (III.2) are such that $\vec{n}_{i}$ are in the same plane of the Bloch sphere, then they are not fully incompatible w.r.t. disjoint-convex-mixing.
 Proof.
-Note that permutation of the outcomes of $M_{i}$ in (12) is nothing but taking $-\vec{n}_{i}$ instead of $\vec{n}_{i}$.
-If the three $\vec{n}_{i}$ are in the same plane, then there exists at least one triple $(i,j,k)$ and permutations of outcomes such that
+If the three $\vec{n}_{i}$ are in the same plane, then there exists at least one triple $(i,j,k)$ such that
 $\vec{n}_{0}=\frac{1}{c}(q\vec{n}_{1}+(1-q)\vec{n}_{2})$ (17)
 for some $c\in(0,1],q\in[0,1]$.
-In other words, there exists a triple $(i,j,k)$ and permutation of outcomes such that $\vec{n}_{i}$ is expressed as a linear combination of $\vec{n}_{j}$ and $\vec{n}_{k}$ with non-negative coefficients (here $q/c$ and $(1-q)/c$), such that sum of those two non-negative coefficients is greater than or equal to 1.
-Substituting this into left hand side of (13) taking $+$ sign, we find
+In other words, there exists a triple $(i,j,k)$ so that $\vec{n}_{i}$ is expressed as a linear combination of $\vec{n}_{j}$ and $\vec{n}_{k}$ with non-negative coefficients (here $q/c$ and $(1-q)/c$), where the sum of those two non-negative coefficients is greater than or equal to 1.
+Substituting this into left hand side of (16), we find
 $||(1+c)\vec{n}_{i}||+||(1-c)\vec{n}_{i}||\leqslant|1+c|+|1-c|=2.$ (18)
 […]
-Observation 4.
-Suppose three qubit measurements (12) are such that $\vec{n}_{0}=\nu_{0}\hat{x}$, $\vec{n}_{1}=\nu_{1}\hat{y}$, $\vec{n}_{2}=\nu_{2}\hat{z}$ with $0\leq\nu_{1},\nu_{2},\nu_{3}\leq 1$, that is, they are the noisy version of Pauli observables,
-$M_{i}:=\frac{1}{2}\left(\mathbbm{1}\pm\nu_{i}\sigma_{i}\right)=\nu_{i}\left(\frac{\mathbbm{1}\pm\sigma_{i}}{2}\right)+(1-\nu_{i})\frac{\mathbbm{1}}{2},$ (16)
-taking $i=0,1,2$ and $(\sigma_{0},\sigma_{1},\sigma_{2})=(\sigma_{x},\sigma_{y},\sigma_{z})$. Then they are fully incompatible w.r.t convex mixture if and only if
+Theorem 3.
+Consider three-qubit measurements (III.2) are such that $\vec{n}_{0}=\nu_{0}\hat{x}$, $\vec{n}_{1}=\nu_{1}\hat{y}$, $\vec{n}_{2}=\nu_{2}\hat{z}$ with $0\leqslant\nu_{0},\nu_{1},\nu_{2}\leqslant 1$, that is, the noisy version of Pauli observables,
+$M_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\nu_{0}\sigma_{x}\right)=\nu_{0}\left(\frac{\mathbbm{1}+(-1)^{z}\sigma_{x}}{2}\right)+(1-\nu_{0})\frac{\mathbbm{1}}{2},$
+$N_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\nu_{1}\sigma_{y}\right)=\nu_{1}\left(\frac{\mathbbm{1}+(-1)^{z}\sigma_{y}}{2}\right)+(1-\nu_{1})\frac{\mathbbm{1}}{2},$
+$R_{z}=\frac{1}{2}\left(\mathbbm{1}+(-1)^{z}\nu_{2}\sigma_{z}\right)=\nu_{2}\left(\frac{\mathbbm{1}+(-1)^{z}\sigma_{z}}{2}\right)+(1-\nu_{2})\frac{\mathbbm{1}}{2},$
+with $z=0,1$. These measurements are fully incompatible w.r.t. disjoint-convex-mixing if and only if
 $\min\bigg{\{}\nu^{2}_{0}+\frac{\nu^{2}_{1}\nu^{2}_{2}}{\nu^{2}_{1}+\nu^{2}_{2}},\nu^{2}_{1}+\frac{\nu^{2}_{0}\nu^{2}_{2}}{\nu^{2}_{0}+\nu^{2}_{2}},\nu^{2}_{2}+\frac{\nu^{2}_{0}\nu^{2}_{1}}{\nu^{2}_{0}+\nu^{2}_{1}}\bigg{\}}>1.$ (20)
@@
 $=2\sqrt{\nu^{2}_{i}+\frac{\nu^{2}_{j}\nu^{2}_{k}}{\nu^{2}_{j}+\nu^{2}_{k}}}$ (21)
-Thus, the r.h.s of (13) is greater than 2 for all values of $q$ if and only if
+Thus, the right-land-side of (16) is greater than 2 for all values of $q$ if and only if
 $\nu^{2}_{i}+\frac{\nu^{2}_{j}\nu^{2}_{k}}{\nu^{2}_{j}+\nu^{2}_{k}}>1.$ (22)
 […]
-Clearly, the three Pauli observables are fully incompatible w.r.t. convex mixture, and moreover, if we take an equal amount of noise $\nu_{0}=\nu_{1}=\nu_{2}=\nu$, then (17) implies $\nu>\sqrt{2/3}$.
-We can generalize this notion for a set of $n$ measurements.
-Definition 4 ($k$-incompatible measurements w.r.t. convex mixtures).
-A set of measurements is $k$-incompatible w.r.t convex mixtures if after taking all possible convex mixtures over all possible $k$ partitions the resulting $k$ number of measurements from the set of $n$ measurements (with relabelling of outcomes) is incompatible.
-Definition 5 (Fully incompatible measurements w.r.t. convex mixtures).
-A set of measurements is fully incompatible w.r.t convex mixtures if it remains incompatible after taking all possible convex mixtures over all possible partitions of $n$ measurements (with relabelling of outcomes).
-Theorem 2.
-Fully incompatible measurements w.r.t convex mixtures imply that every pair of measurements from that set is incompatible. The reverse implication does not hold.
-Consider a set of $n$ measurements, $M=\{M_{1},M_{2},M_{3},\ldots,M_{n}\}$, in which there is a pair of measurements $M_{i}$ and $M_{j}$ which are compatible.
-Now if we make two partitions where $M_{i}$ and $M_{j}$ are in different partitions, the convex mixture of measurements must be compatible.
-This is true for any compatible pair $M_{i}$ and $M_{j}$.
-Thus, if the measurements are fully incompatible w.r.t convex mixture, every pair of measurements must necessarily be incompatible.
-The reverse is not true.
-Consider the three noisy Pauli measurements of Eq.(16).
-It can be shown by using semi-definite programming that if $0.71<\nu\leq 0.81$, the measurements are pairwise incompatible but it is not fully incompatible w.r.t convex mixture.
-III Robustness under noise for different levels of incompatibility
+Clearly, the three Pauli observables are fully incompatible w.r.t.
+disjoint-convex-mixing, and moreover, if we take an equal amount of noise $\nu_{0}=\nu_{1}=\nu_{2}=\nu$, then (20) implies $\nu>\sqrt{2/3}$.
+We can generalize the notion of full incompatibility w.r.t. disjoint-convex-mixing.
+Definition 4 ($k$-incompatible measurements w.r.t. disjoint-convex-mixing).
+Given a set of $n$ measurements, the measurements are $k$-incompatible w.r.t.
+disjoint-convex-mixing if, after taking every possible disjoint-convex-mixing that yields $k$ number of measurements, the resulting measurements are incompatible.
+Definition 5 (Fully incompatible measurements w.r.t. disjoint-convex-mixing).
+A set of $n$ measurements is fully incompatible w.r.t. disjoint-convex-mixing if it is $k$-incompatible for all $k=2,\cdots,n$.
+Observation 3.
+Fully incompatible measurements w.r.t. disjoint-convex-mixing imply that every pair of measurements from that set is incompatible. The reverse implication does not hold.
+Proof.
+Consider a set of $n$ measurements, in which there is a pair of measurements $\{M_{z}\}$ and $\{N_{z}\}$ that are compatible.
+Now if we make two partitions where $\{M_{z}\}$ and $\{N_{z}\}$ are in different partitions and the disjoint-convex-mixing is such that the probabilities of arising all other measurements are zero, then the resultant measurement pair must be compatible.
+This is true for any compatible pair of measurements.
+Thus, if the measurements are fully incompatible w.r.t.
+disjoint-convex-mixing, every pair of measurements must necessarily be incompatible.
+[…]
+Some explicit examples of $k$-incompatible measurements w.r.t. both of these operations are provided in the subsequent section.
+IV Robustness under noise for different levels of incompatibility
 In this section, we analyze the role of noise on quantum measurements and study how the incompatibility properties depend on it.
+More specifically, we study the variation of the critical amount of noise in the different layers of incompatibility under these classical operations (coarse-graining and disjoint-convex-mixing).
 Due to the ubiquitous nature of noise, it is pertinent to study the extent to which noise could be tolerated by a set of measurements while still retaining their incompatibility.
-We take a noisy version of mutually unbiased bases measurements, $M_{i|x}\equiv\{\nu|\phi_{i}^{x}\rangle\langle\phi_{i}^{x}|+\frac{(1-\nu)}{d}\mathbbm{1}_{d}\},\,\,\text{where}\,\,\{|\phi_{i}^{x}\rangle\}$ form mutually unbiased bases measurements in $\mathbb{C}^{d}$.
+We take a noisy version of mutually unbiased bases measurements (MUBs),
+$M_{i|x}=\nu|\phi_{i|x}\rangle\langle\phi_{i|x}|+\frac{(1-\nu)}{d}\mathbbm{1}_{d},$ (23)
+where $\{|\phi_{i|x}\rangle\}_{i,x}$ form mutually unbiased bases measurements in $\mathbb{C}^{d}$.
 Here $\nu$ is the robustness parameter (or visibility parameter) and $(1-\nu)$ is the noise parameter, $0\leq\nu\leq 1$.
 When the noise parameter is zero (i.e., robustness, $\nu=1$), the measurements are fully incompatible, and when the noise parameter is one (i.e., the robustness, $\nu=0$), the measurements are trivial and compatible.
-Our aim is to obtain the critical value of the robustness parameter above which the measurements remain incompatible.
+Our aim is to obtain the critical value of the robustness parameter above which the measurements remain incompatible before and after different classical operations.
 To check the compatibility, i.e., the existence of a parent POVM, we use the method described in Gühne et al.
 (2023).
-This can be casted as a semi-definite programming (SDP) problem that takes a set of measurements $\{M_{z_{x}|x}\}$ and deterministic classical post-processings $p(z_{x}|x,\lambda)$ as input, and checks whether the measurements are compatible or not, subject to the constraints
+This can be cast as a semi-definite programming (SDP) problem that takes a set of measurements $\{M_{z_{x}|x}\}$ and deterministic classical post-processings $p(z_{x}|x,\lambda)$ as input, and checks whether the measurements are compatible or not, subject to the constraints
 $\sum_{\lambda}p(z_{x}|x,\lambda)G_{\lambda}=M_{z_{x}|x}\forall x,z_{x},$ (24)
@@
 This method finds the maximum value of $\mu$ for each $\{p(z_{x}|x,\lambda)\}$.
-If this optimization returns a negative value of $\mu$, then the constraint of Eq.(22) can’t be fulfilled, which implies that the measurements $\{M_{z_{x}|x,\lambda}\}$ are incompatible.
+If this optimization returns a negative value of $\mu$, then the constraint of Eq.(26) cannot be fulfilled, which implies that the measurements $\{M_{z_{x}|x}\}$ are incompatible.
 Otherwise, they are compatible.
-III.1 Coarse-graining of outcomes
-Dimension $3$. Let $\{P_{i}\}$ and $\{Q_{j}\},i,j\in[3]$ be two measurements acting on $\mathbb{C}^{3}$, where
-$P_{i}=\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{3},$
-$Q_{j}=\nu|\psi_{j}\rangle\langle\psi_{j}|+(1-\nu)\frac{\mathbbm{1}}{3},$ (23)
+As discussed earlier, the degree of incompatibility of measurements may be reduced due to coarse-graining of outcomes and convex mixing of measurements.
+In the subsequent subsection, we present an explicit analysis of how the degree of incompatibility (quantified by the robustness parameter (Heinosaari et al., 2015)) varies as the measurements are subjected to coarse-graining and disjoint-convex-mixing, in dimensions $3$ and $4$.
+For simplicity, we have considered the measurements to be MUBs (as discussed earlier), where the number of outcomes are same as that of their dimension.
+Also, as MUBs exhibit the highest degree of incompatibility, (Designolle et al., 2019b) it is a natural choice to use them for the prominent observation of the various layers of incompatibility under the classical operations.
+However, one can apply this formalism to any set of measurements in a similar fashion.
+IV.1 Robustness of incompatibility under coarse-graining of outcomes
+Dimension $3$. Let $\{M_{i}\}$ and $\{N_{j}\},i,j\in\{1,2,3\}$ be two three-outcome measurements acting on $\mathbb{C}^{3}$, where
+$M_{i}=\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{3},$
+$N_{j}=\nu|\psi_{j}\rangle\langle\psi_{j}|+(1-\nu)\frac{\mathbbm{1}}{3},$ (27)
 with
@@
 $|\psi_{1}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega|1\rangle+\omega^{2}|2\rangle),$
-$|\psi_{2}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega^{2}|1\rangle+\omega|2\rangle),$
-It is observed that for $\nu>0.683$, the two measurements are incompatible.
-For finding ”$2$-incompatible w.r.t coarse-graining”, we consider all possible coarse-graining of outcomes of the two measurements such that after coarse-graining, the new set of measurements have at least two outcomes each.
-We find that these measurements are incompatible w.r.t all possible coarse-graining for $\nu>0.711$.
-Note that for $0.683<\nu\leq 0.711$, the incompatible measurements become compatible under certain coarse-graining of outcomes (see TABLE 1).
+$|\psi_{2}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega^{2}|1\rangle+\omega|2\rangle),$ (28)
+$\omega$ being the cube roots of unity.
+It can be easily checked by SDP that the measurements in Eqs.(27) and (28) are incompatible (or equivalently, $3$-incompatible using the definition 2 ) for robustness parameter $(\nu)>0.683$.
+Now, any non-trivial coarse-graining reduces the measurement outcomes to two, so one can check the incompatibility after all possible choices of such non-trivial coarse-grainings.
+The measurements are $2$-incompatible (by the definition 2) for $\nu>0.711$.
+For $0.683<\nu\leq 0.711$, the measurements are $3$-incompatible, but not $2$-incompatible.
+This represents the case where the initial measurements, although taken to be incompatible, become compatible after a non-trivial coarse-graining.
+However, for $\nu>0.711$, the measurements are $2$-incompatible (and hence $3$-incompatible), suggesting that they remain incompatible after all non-trivial coarse-grainings.
+These findings suggest that $2$-incompatible measurements exhibit a stronger form of incompatibility compared to the $3$-incompatible measurements (hence supporting Observation 1), and hence are potential candidates for any resource-theoretic applications that harness incompatibility.
+Note that for three outcome measurements, any non-trivial coarse-graining will reduce the outcomes to two.
+So, the number of outcomes is always $\leq 3$ (where $3$ corresponds to the case where there is no coarse-graining).
+So, $4$-incompatibility is not applicable (N.A.) here.
 Dimension $4$.
-For checking incompatibility w.rt coarse-graining in $\mathbb{C}^{4}$, the same procedure is repeated for two POVM measurements with $4$ outcomes each.
-The corresponding measurements are $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\},i,j\in[4]$.
-$P^{\prime}_{i}$ $=$ $\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{4},$
-$Q^{\prime}_{j}$ $=$ $\nu|\psi^{\prime}_{j}\rangle\langle\psi^{\prime}_{j}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (24)
+For checking incompatibility w.r.t coarse-graining in $\mathbb{C}^{4}$, the same procedure is repeated taking two POVM measurements with four outcomes each.
+The corresponding measurements are $\{M^{\prime}_{i}\},\{N^{\prime}_{j}\},i,j\in\{1,2,3,4\}$:
+$M^{\prime}_{i}$ $=$ $\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{4},$
+$N^{\prime}_{j}$ $=$ $\nu|\psi^{\prime}_{j}\rangle\langle\psi^{\prime}_{j}|+(1-\nu)\frac{\mathbbm{1}}{4},$ (29)
 with
@@
 $|\psi^{\prime}_{2}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle-|2\rangle+|3\rangle),$
-$|\psi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+|2\rangle-|3\rangle),$ (25)
-for $\nu>0.666$, these measurements are incompatible.
-We find that for $\nu>0.675$, the measurements are incompatible under all possible non-trivial coarse-graining that gives rise to at least $3$ outcomes.
-The measurements are fully incompatible for $\nu>0.72$.
-The results are summarized in TABLE 1.
+$|\psi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+|2\rangle-|3\rangle).$ (30)
+These measurements in Eqs.(IV.1) and (IV.1) are incompatible (or equivalently $4$-incompatible, using the definition 2) for $\nu>0.666$.
+In this case, any non-trivial coarse-graining will reduce the measurement outcomes to three or two, and hence we shall have $3$-incompatiblity and $2$-incompatiblity respectively.
+For $0.666<\nu\leq 0.675$, the incompatible measurements become compatible after all possible coarse-grainings that reduce the outcomes to at least three.
+Further reducing the number of outcomes by coarse-graining gives a tighter bound for the critical value of the robustness parameter above which the measurements are incompatible.
+Hence, for $0.675<\nu\leq 0.720$, the measurements are $3$-incompatible but become compatible if one further reduces the number of outcomes by coarse-graining.
+These measurements exhibit the strongest form of incompatibility for $\nu>0.720$ (as also discussed in Observation 1), i.e., they remain incompatible after all possible non-trivial coarse-grainings.
+The results for coarse-graining of measurement outcomes in dimensions $3$ and $4$ are summarized in TABLE 1.
 In dimension $3$ In dimension $4$
-Incompatible 0.683 0.666
+$4$-incompatible N.A. 0.666
+$3$-incompatible 0.683 0.675
 $2$-incompatible 0.711 0.720
-3-incompatible 0.683 0.675
-Table 1: Critical values of robustness for MUBs w.r.t coarse-graining in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
-III.2 Convex mixing of measurements
+Table 1: Critical values of robustness for MUBs w.r.t. coarse-graining in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$, above which the measurements are incompatible.
+IV.2 Robustness of incompatibility under disjoint-convex-mixing of measurements
 Dimension $3$.
-For checking incompatibility w.r.t convex mixtures (CM), a minimum of three measurements are needed.
-The two measurements are $\{P_{i}\}$ and $\{Q_{j}\}$, as given in Eqs.(23), The third measurement is $\{R_{k}\},k\in[3]$ where
+For checking incompatibility w.r.t.
+disjoint-convex-mixing (CM), a minimum of three measurements are needed.
+The two measurements are $\{M_{i}\}$ and $\{N_{j}\}$, as given in Eqs.(27) and Eq.(28).
+The third measurement is $\{R_{k}\},k\in\{1,2,3\}$ where
 $R_{k}=\nu|\phi_{k}\rangle\langle\phi_{k}|+(1-\nu)\frac{\mathbbm{1}}{3}.$ (31)
 with
-$\begin{split}&|\phi_{0}\rangle=\frac{\omega|0\rangle+|1\rangle+|2\rangle}{\sqrt{3}},|\phi_{1}\rangle=\frac{|0\rangle+\omega|1\rangle+|2\rangle}{\sqrt{3}},\\ &|\phi_{2}\rangle=\frac{|0\rangle+|1\rangle+\omega|2\rangle}{\sqrt{3}},\end{split}$ (27)
-$\omega$ being the cube roots of unity.
-The results are summarized in TABLE 2.
-Dimension $4$. Consider the measurements $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\}$ as defined in Eqs.(III.1) and(III.1) and another measurement $\{R^{\prime}_{k}\}$ where $i,j,k\in[4]$.
-$\{R^{\prime}_{k}\}=\nu|\phi^{\prime}_{k}\rangle\langle\phi^{\prime}_{k}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (28)
+$\begin{split}&|\phi_{0}\rangle=\frac{\omega|0\rangle+|1\rangle+|2\rangle}{\sqrt{3}},\ |\phi_{1}\rangle=\frac{|0\rangle+\omega|1\rangle+|2\rangle}{\sqrt{3}},\\ &|\phi_{2}\rangle=\frac{|0\rangle+|1\rangle+\omega|2\rangle}{\sqrt{3}}.\end{split}$ (32)
+These three measurements are incompatible (equivalently $3$-incompatible, by the definition 4) for $\nu>0.537$.
+From these three incompatible measurements, one can effectively reduce the number of measurements to two, by taking a new measurement as the disjoint-convex-mixing of any two measurements from the initial set of three measurements (this can be done in $3$ ways), while keeping the third measurement same.
+This process of disjoint-convex-mixing may reduce the degree of incompatibility, as discussed before and evidenced for $0.537<\nu\leq 0.764$.
+This corresponds to the case where the two measurements become compatible after considering all possible disjoint-convex-mixing across all partitions.
+For $\nu>0.764$, the measurements are incompatible, suggesting that this is the critical value above which, the measurements are robust against disjoint-convex-mixing.
+Dimension $4$. Consider the measurements $\{M^{\prime}_{i}\},\{N^{\prime}_{j}\}$ as defined in Eqs. (IV.1) and (IV.1) and another measurement $\{R^{\prime}_{k}\}$ where $i,j,k\in\{1,2,3,4\}$,
+$R^{\prime}_{k}=\nu|\phi^{\prime}_{k}\rangle\langle\phi^{\prime}_{k}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (33)
 with
@@
 $|\phi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle+|1\rangle-\mathbbm{i}|2\rangle+\mathbbm{i}|3\rangle).$ (34)
-The results are summarized in TABLE 2.
+These three measurements in dimension $4$ are incompatible for $\nu>0.692$.
+As discussed before, disjoint-convex-mixing makes them compatible and they continue to remain so till $\nu=0.705$.
+So, for $0.692<\nu\leq 0.705$, the three measurements are incompatible, but disjoint-convex-mixing of any two of them makes them compatible.
+For $\nu>0.705$, the measurements are incompatible and robust against disjoint-convex-mixing across any partition that effectively brings down the number of measurements to two.
+The results for disjoint-convex-mixing of measurements in dimensions $3$ and $4$ are summarized in TABLE 2.
 In dimension $3$ In dimension $4$
-Incompatible 0.537 0.692
+$3$-incompatible 0.537 0.692
 $2$-incompatible 0.764 0.705
-3-incompatible 0.537 0.692
-Table 2: Critical values of robustness for MUBs w.r.t convex-mixing in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
-IV Operational witnesses of incompatibility
-IV.1 Coarse-graining of outcomes
-As we mentioned earlier, measurements that are fully incompatible w.r.t coarse-graining show stronger incompatibility compared to the measurements that are not fully incompatible w.r.t coarse-graining.
-To operationally certify whether a set of measurements are fully incompatible w.r.t coarse-graining is important from the perspective of determing the practical utility of such a set for revealing phenomena such as Bell inequality, steering and contextuality Fine (1982a); Wolf et al.
-(2009); Uola et al.
+Table 2: Critical values of robustness for MUBs w.r.t. disjoint-convex-mixing in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$, above which the measurements are incompatible.
+V Operational witnesses of incompatibility
+In this section, we explore how different levels of incompatibility can be witnessed through the input-output statistics derived from characterized devices.
+V.1 Coarse-graining of outcomes
+As we mentioned earlier, measurements that are fully incompatible w.r.t.
+coarse-graining show stronger incompatibility compared to the measurements that are not fully incompatible w.r.t.
+coarse-graining.
+To operationally certify whether a set of measurements is fully incompatible w.r.t.
+coarse-graining is important from the perspective of determining the practical utility of such a set for revealing phenomena such as Bell inequality, steering and contextuality Fine (1982a); Wolf et al.
+(2009a); Uola et al.
 (2014); Quintino et al.
@@
 (2023).
-Operational certification means that we need to infer the incompatibility of the measurements from the input-output statistics of the measurement device without knowing its internal functioning.
-Below we study the two classes of operational certification separately
+Certification means that we need to infer the incompatibility of the measurements from the input-output statistics of the measurement device without knowing its internal functioning.
+Below we study the two classes of certification separately.
 Device-independent witness
-Here one requires neither any prior knowledge of the internal functioning of the measurement device, nor any idea of the dimension of the system on which the measurements act.
-In Wolf et al.
-(2009), it has been proven that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality.
-We can use this result to witness incompatibility for any two 3-outcome projective measurements before and after coarse-graining, from the input-output statistics.
-Semi-Device-independent witness
-Here, although we do not have any prior knowledge of the internal functioning of the measurement device, we do know the dimension of the system on which the measurements act.
-Any two incompatible rank-one projective measurements provide an advantage in $(2,d,d)-$RAC Saha et al.
+Violation of Bell inequality provides device-independent witness for incompatible measurements.
+Here, one neither requires any prior knowledge of the internal functioning of the measurement device nor any idea of the dimension of the system on which the measurements act.
+However, not all incompatible measurements yield violations of Bell inequalities Bene and Vértesi (2018); Hirsch et al.
 (2018).
-If any two POVMs each with $\overline{d}$ outcomes acting on $\mathbb{C}^{d}$ are jointly measurable, the average maximum success probability in $(2,\overline{d},d)$ RAC is bounded by Saha et al.
-(2023)
-$P(2,\overline{d},d)\leq\frac{1}{2}(1+\frac{d}{\overline{d}^{2}})=P_{CB}(2,\overline{d},d),$ (30)
-where $P_{CB}(2,\overline{d},d)$ is the maximum average success probability obtained in a $(2,\overline{d},d)$ RAC using two compatible measurements.
-For full incompatibility of two 3-outcome rank-one projective measurements, we prove the following:
-Theorem 3.
-Two rank-one projective measurements, $M=\Big{\{}|\phi_{0}\rangle,|\phi_{1}\rangle,|\phi_{2}\rangle\Big{\}}$ and $N=\Big{\{}|\psi_{0}\rangle,|\psi_{1}\rangle,|\psi_{2}\rangle\Big{\}}$, can be witnessed to be fully incompatible w.r.t.
-coarse-graining via RAC iff $20<|\langle\phi_{i}|\psi_{j}\rangle|<\frac{4}{5},\quad\forall\,\,i,j=0,1,$ 2 .
+Nonetheless, we have the following results.
+Observation 4.
+Full-incompatibility w.r.t. coarse-graining of any two measurements can always be witnessed in a device-independent way.
 Proof.
-Without loss of generality, any pair of three outcome rank-one projective measurements can be written up to unitary freedom as,
-$M=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|,|2\rangle\langle 2|\Big{\}}\equiv\Big{\{}|\phi_{i}\rangle\langle\phi_{i}|\Big{\}},$
-and
-$N=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|,|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}\equiv\Big{\{}|\psi_{j}\rangle\langle\psi_{j}|\Big{\}},$ (31)
-where $|\psi_{j}\rangle=\alpha_{j}|0\rangle+\beta_{j}|1\rangle+\gamma_{j}|2\rangle$.
-Let us first coarse grain the second and third outcomes. So, the new measurement pair becomes:
-$M^{(2,3)}=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|+|2\rangle\langle 2|\Big{\}},$ (32)
-$N^{(2,3)}=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}.$
-The new compatibility bound in this case is,
 […]
-and the bound is tight. $M^{(2,3)}\,\,\text{and}\,\,N^{(2,3)}$ will give advantage in the (2,2,3) RAC if:
-$\sum_{x,y=0}^{1}$ $||M^{(2,3)}(x)+N^{(2,3)}(y)||>7,\,\,\text{which implies}$
-$||A||$ $+$ $||B||+||C||+||D||>7$ (33)
-$\text{with}\,\,A$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(0)$
-$=$ $|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
-Now, $|\psi_{1}\rangle$ can be written as:
-$|\psi_{1}\rangle=\langle 0|\psi_{1}\rangle|0\rangle+\langle u|\psi_{1}\rangle|u\rangle\,\,\text{with}$ (34)
-$\langle 0|u\rangle=0\,\,\text{and}\,\,|\langle 0|\psi_{1}\rangle|^{2}+|\langle u|\psi_{1}\rangle|^{2}=1.$
-$A$ can be written as a $(2\times 2)$ matrix in the basis $\{|0\rangle,|u\rangle\},$
-$A=\begin{pmatrix}1+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (37)
-So, the maximum eigenvalue of $A\,\,\text{is}\,\,1+|\langle 0|\psi_{1}\rangle|.$ Thus,
-$||A||=1+|\langle 0|\psi_{1}\rangle|.$
-By Eq.(IV.1) we have
-$B$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(1)$ (38)
-$=$ $|0\rangle\langle 0|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|$
-$=$ $\mathbbm{1}+|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|.$
-Similarly, $B$ can be expressed in a block diagonal matrix in the ortho-normal basis $\{|0\rangle,|u\rangle,|v\rangle\},$
-$B=\begin{pmatrix}\Gamma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Gamma=\begin{pmatrix}2+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (43)
-and
-$||B||=2+|\langle 0|\psi_{1}\rangle|.$
-Also, from Eq.(IV.1)
-$C$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(0)$ (44)
-$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{1}\rangle\langle\psi_{1}|\,\,$
-$=$ $\mathbbm{1}-|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
-Similarly, $C$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
-$C=\begin{pmatrix}\Sigma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Sigma=\begin{pmatrix}|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (49)
-and
-$||C||=1+\sqrt{1-|\langle 0|\psi_{1}\rangle|^{2}}.$
-Similarly, from Eq.(IV.1)
-$D$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(1)$ (50)
-$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\,\,$
-$=$ $2\mathbbm{1}-|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$
-and $D$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
-$D=\begin{pmatrix}\Xi&0\\ 0&2\end{pmatrix}\,\,\text{where}\,\,\Xi=\begin{pmatrix}1-|\langle 0|\psi_{1}\rangle|^{2}&-\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ -\langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&2-|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (55)
-and $||D||=2.$
-[…]
-Device-independent witness of full-incompatibility of two measurements w.r.t coarse-graining from a single experiment
-We now study the following problem: given two measurements of outcomes $d_{1}$ and $d_{2}$ respectively, how can one witness the full-incompatibility w.r.t coarse-graining in a device-independent way from a single experiment?
-Here ”single experiment” implies the Bell test which gives a Bell-inequality violation for the two measurements of outcomes $d_{1}$ and $d_{2}$ respectively as chosen as inputs of Alice and some measurement choices for Bob and a shared entangled state.
-As we know if two measurements are $2-$ incompatible w.r.t coarse-graining then they are fully-incompatible w.r.t coarse-graining.
-Also, we use the result of Wolf et al.
-(2009), that any two binary outcome incompatible measurements violate the Bell-CHSH inequality.
-Thus, from the original measurement statistics we do all possible coarse-grainings to make it two outcome measurements and check the Bell-CHSH inequalities.
-If, all possible coarse-grained two-outcome measurements violate the Bell-CHSH inequalities then the measurements are fully-incompatible w.r.t coarse-graining.
-To understand how this technique works, we consider a well-known example of two three-outcome rank-one projective measurements in $\mathbb{C}^{3}$ which give maximum violation of the CGLMP inequality Collins et al.
+Device-independent witness from a single Bell experiment
+In the above-mentioned approach, different Bell experiments are probed for different coarse-graining.
+It is interesting to explore whether different levels of incompatibility w.r.t.
+coarse-graining can be witnessed universally through a single Bell experiment in a device-independent manner.
+In this case, the objective is to consider different coarse-grained statistics of a Bell experiment and check whether those statistics have local explanations or not.
+To understand how this technique works, we consider a well-known example of two three-outcome rank-one projective measurements in $\mathbb{C}^{3}$ which give maximum violation of the Collins-Gisin-Linden-Massar-Popescu (CGLMP) inequality Collins et al.
 (2002), a well-studied Zohren and Gill (2008); Das et al.
@@
 with, $b\in\{1,2\}$ corresponding to two different measurement settings of Bob, and $\eta\in\{0,1,2\},\beta_{1}=\frac{1}{4},\beta_{2}=-\frac{1}{4}.$
+Specifically, here we investigate the different layers of incompatibility of the measurements given in (35) under coarse-graining.
 The incompatibility of the measurements of Eq.(35) before coarse-graining is guaranteed by the CGLMP-inequality violation.
-Now we investigate the incompatibility status of the measurements of Alice after all possible coarse-graining from the measurement statistics obtained in the CGLMP experiment.
-Alice’s measurements now have two outcomes each and it is a (2,2,2,3) scenario, i.e., two inputs for Alice and two inputs for Bob and each input of Alice has two outcomes and each input of Bob has three outcomes.
+Now we give a theoretical analysis to investigate the incompatibility status of the measurements of Alice given in (35) after all possible coarse-grainings.
+The results are summarized in TABLE   3.
+Theoretical explanation.— Out of three outcomes if any of the two outcomes are coarse-grained it becomes a (2,2,2,3) scenario, i.e., two inputs for Alice and two inputs for Bob and each input of Alice has two outcomes and each input of Bob has three outcomes.
 For this scenario, Collins and Gisin have shown that there are a total of 72 CH-facet inequalities Collins and Gisin (2004).
 The CH-inequalities Fine (1982b) are of the form:
-$-1\leq$ $P(00|A_{1},B_{1})+P(00|A_{1},B_{2})+P(00|A_{2},B_{2})$ (58)
-$-P(00|A_{2},B_{1})-P(0|A_{1})-P(0|B_{2})\leq 0.$
+$-1\leq S\leq 1,$ (37)
+with CH-functional
+$S=$ $P(00|A_{1},B_{1})+P(00|A_{1},B_{2})+P(00|A_{2},B_{2})$ (38)
+$-P(00|A_{2},B_{1})-P(0|A_{1})-P(0|B_{2}).$
 The other CH-inequalities are obtained by (1) interchanging $A_{1}$ with $A_{2}$, (2) interchanging $B_{1}$ with $B_{2}$, and (3) interchanging both $A_{1}$ with $A_{2}$ and $B_{1}$ with $B_{2}$.
@@
 $P(\overline{0}|A_{i})=P(0|A_{i})+P(1|A_{i}),$
-$P(\overline{0}|B_{i})=P(0|B_{i})+P(1|B_{i}),\,\,i,j\in\{1,2\}.$ (61)
-Similarly, there are other 8 possible clubbings for Bob’s measurement outcomes and for each clubbing, we have facet inequalities similar to Eq.(59).
+$P(\overline{0}|B_{j})=P(0|B_{j})+P(1|B_{j}),\,\,i,j\in\{1,2\}.$ (41)
+Similarly, there are eight other possible clubbings for Bob’s measurement outcomes and for each clubbing, we have facet inequalities similar to Eq.(39).
 One can check that when there are the same coarse-graining of outcomes for both measurement inputs of Alice, we get a CH-inequality violation.
@@
 CH-violation under a particular coarse-graining signifies that the measurements are incompatible.
-However, when there is no CH-violation we can not conclude anything regarding the incompatibility status.
+However, when there is no CH-violation, we can not conclude anything regarding the incompatibility status.
 This is depicted in TABLE   3.
-$A_{1}$ $A_{2}$ Incompatible
-(0,1) (0,1) yes
-(0,1) (1,2) ?
-(0,1) (0,2) yes
-(1,2) (0,1) yes
-(1,2) (1,2) yes
-(1,2) (0,2) ?
-(0,2) (0,1) ?
-(0,2) (1,2) yes
-(0,2) (0,2) yes
-Table 3: The incompatibility status for all possible coarse-graining (CG) of outcomes for Alice’s measurements ($A_{1}$ and $A_{2}$) as inferred from the statistics of CGLMP experiment.
-Here “yes” denotes that under the particular CG of $A_{1}$ and $A_{2}$, the measurement remains incompatible.
-On the other hand, “????” denotes that we can not make any conclusion about their incompatibility.
-IV.2 Convex mixing of measurements
+Possible
+CG
+choices $\Delta S_{T}$ $\Delta S_{E}$ Incompatible
+of
+$A_{1}$ $A_{2}$
+(0,1) (0,1) 0.126 0.122 yes
+(0,1) (1,2) 0 0 ?
+(0,1) (0,2) 0.126 0.122 yes
+(1,2) (0,1) 0.126 0.122 yes
+(1,2) (1,2) 0.126 0.122 yes
+(1,2) (0,2) 0 0 ?
+(0,2) (0,1) 0 0 ?
+(0,2) (1,2) 0.126 0.122 yes
+(0,2) (0,2) 0.126 0.122 yes
+Table 3: Incompatibility status for all possible coarse-graining (CG) choices of outcomes of Alice’s measurements ($A_{1}$ and $A_{2}$) in the CGLMP scenario.
+Here $\Delta S_{T}$ refers to the amount of CH-inequality violation obtained in the theory and $\Delta S_{E}$ refers to the CH-inequality violation calculated from the states and measurements viz.
+(42) and (43) realised in the experiment Zhang et al.
+(2024).
+Here “yes” denotes that under the particular CG of $A_{1}$ and $A_{2}$, the measurements can be witnessed to be incompatible as they give CH-inequality violation.
+On the other hand, “????” denotes that we can not make any conclusion about their incompatibility under those particular CG as there is no violation of CH-inequality in those cases.
+Experimental realization.— Recently, an experimental study on CGLMP inequality has been performed Zhang et al.
+(2024) where the two parties share orbital angular momentum entanglement in a scenario of multiple settings and outcomes.
+In their experiment, for $d=3$ they prepared an orbital angular momentum entangled state which is in $\mathbb{C}^{3}\otimes\mathbb{C}^{3}$ of the form:
+$|\psi\rangle$ $=$ $0.596|+1\rangle_{A}|-1\rangle_{B}+0.529|+2\rangle_{A}|-2\rangle_{B}$ (42)
+$+0.604|-1\rangle_{A}|+1\rangle_{B},$
+with $|+l\rangle_{A}$ corresponds to the orbital angular momentum eigenstate of the signal photon (A) with orbital angular momentum $+l\hbar\,\,\text{and}\,\,|-l\rangle_{B}$ corresponds the orbital angular momentum eigenstate of the corresponding idler photon (B) with orbital angular momentum $-l\hbar.$ The von Neuman measurements of Alice and Bob are given by $A_{i}\equiv\{|\Gamma^{i}_{s}\rangle_{A}\langle\Gamma^{i}_{s}|\}\,\,\text{and}\,\,B_{j}\equiv\{|\Theta^{j}_{t}\rangle_{B}\langle\Theta^{j}_{t}|\}$ respectively, where
+$|\Gamma^{i}_{s}\rangle_{A}=\frac{1}{\sqrt{3}}(|+1\rangle_{A}+\omega^{s+\sigma_{i}}|+2\rangle_{A}+\omega^{2(s+\sigma_{i})}|-1\rangle_{A}),$ (43)
+and
+$|\Theta^{j}_{t}\rangle_{B}=\frac{1}{\sqrt{3}}(|-1\rangle_{B}+\omega^{-t-\gamma_{j}}|-2\rangle_{B}+\omega^{2(-t-\gamma_{j})}|+1\rangle_{B}),$ (44)
+with $i,j\in\{1,2\}\,\,\text{and}\,\,s,t\in\{0,1,2\},\sigma_{1}=\frac{1}{4},\sigma_{2}=\frac{3}{4},\gamma_{1}=\frac{1}{2}\,\,\text{and}\,\,\gamma_{2}=0,\,\,\omega=\text{exp}(\mathbbm{1}\frac{2\pi}{3}).$ For $d=3$ the incompatibility status of the measurements (see (43)) used in their experimental arrangements before and after coarse-graining is depicted in TABLE 3.
+Semi-device-independent witness
+In the semi-device-independent approach in prepare-and-measure experiments, we do not have any prior knowledge of the internal functioning of the measurement device; however, we assume the dimension of the system on which the measurements act.
+To witness different levels of incompatibility, here we focus on a class of communication tasks, namely, $(2,\overline{d},d)-$ random access code tasks (RAC).
+In this task, the sender, Alice, gets two-dit string input message $(x_{1},x_{2})$ with $x_{1},x_{2}\in[\overline{d}],$ and can communicate a $d-$dimensional system to the receiver, Bob, who wants to guess the value of any of the two dits, i.e.
+$x_{1}$ or $x_{2}$, randomly.
+The notation of $(2,\overline{d},d)-$RAC is adopted from Saha et al.
+(2023).
+If any two POVMs, each with $\overline{d}$ outcomes acting on $\mathbb{C}^{d}$ are jointly measurable, the average success probability of this task,
+$P(2,\overline{d},d)\leq P_{CB}(2,\overline{d},d)=\frac{1}{2}\left(1+\frac{d}{\overline{d}^{2}}\right),$ (45)
+where $P_{CB}(2,\overline{d},d)$ is an upper bound on average success probability using two compatible measurements Carmeli et al.
+(2020); Saha et al.
+(2023).
+It turns out any two incompatible rank-one projective measurements can be witnessed through $(2,d,d)-$RAC Saha et al.
+(2020).
+We prove the following result for full incompatibility of two 3-outcome rank-one projective measurements w.r.t.
+coarse-graining.
+Theorem 4.
+Two 3-outcome rank-one projective measurements, $M=\left\{|\phi_{0}\rangle,|\phi_{1}\rangle,|\phi_{2}\rangle\right\}$ and $N=\left\{|\psi_{0}\rangle,|\psi_{1}\rangle,|\psi_{2}\rangle\right\}$, can be witnessed to be fully incompatible w.r.t.
+coarse-graining via RAC if and only if $20<|\langle\phi_{i}|\psi_{j}\rangle|<\frac{4}{5},\quad\forall\,\,i,j=0,1,$ 2 .
+Proof.
+Without loss of generality, any pair of three outcome rank-one projective measurements can be written up to unitary freedom as,
+$M=\left\{|0\rangle\langle 0|,|1\rangle\langle 1|,|2\rangle\langle 2|\right\}\equiv\left\{|\phi_{i}\rangle\langle\phi_{i}|\right\},$
+$N=\left\{|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|,|\psi_{3}\rangle\langle\psi_{3}|\right\},$ (46)
+where $|\psi_{j}\rangle=\alpha_{j}|0\rangle+\beta_{j}|1\rangle+\gamma_{j}|2\rangle$.
+Let us first consider the coarse-graining of the second and third outcomes for both measurements. So, the new measurement pair becomes:
+$M^{(2,3)}=\big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|+|2\rangle\langle 2|\big{\}},$ (47)
+$N^{(2,3)}=\big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\big{\}}.$
+This pair of measurements appears in $(2,2,3)-$RAC and the compatibility bound in this case is,
+$P_{CB}(2,2,3)=\frac{1}{2}\left(1+\frac{3}{2^{2}}\right)=\frac{7}{8}$ (48)
+by Eq.(45) and this bound is tight. $M^{(2,3)}\,\,\text{and}\,\,N^{(2,3)}$ will give advantage in $(2,2,3)-$RAC if
+$\sum_{x,y=0}^{1}||M^{(2,3)}_{x}+N^{(2,3)}_{y}||=\sum_{i=1}^{4}||A_{i}||>7,$ (49)
+with
+$A_{1}=M^{(2,3)}_{0}+N^{(2,3)}_{0}=|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|,$
+$A_{2}=M^{(2,3)}_{0}+N^{(2,3)}_{1}=\mathbbm{1}+|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$
+$A_{3}=M^{(2,3)}_{1}+N^{(2,3)}_{0}=\mathbbm{1}-|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|,$
+$A_{4}=M^{(2,3)}_{1}+N^{(2,3)}_{1}=2\mathbbm{1}-|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$ (50)
+and ||.||||.||| | . | | denotes the maximum eigenvalue of an operator. Now, $|\psi_{1}\rangle$ can be written as follows
+$|\psi_{1}\rangle=\langle 0|\psi_{1}\rangle|0\rangle+\langle u|\psi_{1}\rangle|u\rangle\,\,\text{with}$ (51)
+$\langle 0|u\rangle=0\,\,\text{and}\,\,|\langle 0|\psi_{1}\rangle|^{2}+|\langle u|\psi_{1}\rangle|^{2}=1,$
+for some vector $|u\rangle.$ So, $A_{1}$ can be written as a $(2\times 2)$ matrix in the basis $\{|0\rangle,|u\rangle\},$
+$A_{1}=\begin{pmatrix}1+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (52)
+the maximum eigenvalue of $A_{1}\,\,\text{is}\,\,1+|\langle 0|\psi_{1}\rangle|.$ Thus,
+$||A_{1}||=1+|\langle 0|\psi_{1}\rangle|.$ (53)
+Similarly, $A_{2}$ can be expressed in a block diagonal matrix in the ortho-normal basis $\{|0\rangle,|u\rangle,|v\rangle\},$
+$A_{2}=\begin{pmatrix}\Gamma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Gamma=\begin{pmatrix}2+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (54)
+and
+$||A_{2}||=2+|\langle 0|\psi_{1}\rangle|.$ (55)
+$A_{3}$ and $A_{4}$ can also be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
+$A_{3}=\begin{pmatrix}\Sigma&0\\ 0&1\end{pmatrix},\ \Sigma=\begin{pmatrix}|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (56)
+and
+$||A_{3}||=1+\sqrt{1-|\langle 0|\psi_{1}\rangle|^{2}}.$ (57)
+$A_{4}=\begin{pmatrix}\Xi&0\\ 0&2\end{pmatrix},\ \Xi=\begin{pmatrix}1-|\langle 0|\psi_{1}\rangle|^{2}&-\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ -\langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&2-|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (58)
+and $||A_{4}||=2.$
+[…]
+V.2 Disjoint-convex-mixing of measurements
+We now study the operational witness of incompatibility w.r.t. disjoint-convex-mixing of measurements.
 Device-independent witness
-Here, one can simply use the result that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality Wolf et al.
-(2009).
-Considering the convex mixture of measurements for any three binary-outcome measurements, this result can be used to witness full incompatibility w.r.t.
-convex mixing of measurements.
+Given a set of $n$ measurements, we can make $k$ partitions of it, where $k\in$$\{2,\cdots,n\}$.
+If the $k$ number of measurements obtained by the disjoint-convex-mixing of the measurements provides a violation of Bell inequalities for all possible disjoint-convex-mixing and permutations, then the measurements are witnessed to be $k-$incompatible w.r.t.
+disjoint-convex-mixing in a device-independent way.
+Using the result that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality Wolf et al.
+(2009a), we can witness 2-incompatibility w.r.t.
+disjoint-convex-mixing from any set of binary-outcome measurements.
 Semi-device-independent witness
-For three or more numbers of measurements we can witness whether the measurements are fully-incompatible w.r.t convex-mixture in a semi-device independent manner by constructing a suitable random access codes task.
+For three or more numbers of measurements, we can witness whether the measurements are fully-incompatible w.r.t.
+disjoint-convex-mixture in a semi-device independent manner by constructing suitable random access code tasks.
 Theorem 5.
-Three noisy Pauli measurements of Eq.(16) are witnessed to be fully incompatible w.r.t convex-mixing via RAC iff: $\sqrt{\frac{2}{3}}<\nu\leq 1.$
+Three noisy Pauli measurements of Eq.
+(3) with equal noise ($\nu=\nu_{0}=\nu_{1}=\nu_{2}$) are witnessed to be fully incompatible w.r.t.
+disjoint-convex-mixing via RAC if and only if $\sqrt{2/3}<\nu\leqslant 1.$
 Proof.
-Let us denote the measurements of Eq.(16) as $\mathcal{M}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{0}\sigma_{x}\right)$, $\mathcal{M}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{0}\sigma_{x}\right)$, $\mathcal{N}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{1}\sigma_{y}\right)$, $\mathcal{N}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{1}\sigma_{y}\right)$, $\mathcal{R}_{0}=\frac{1}{2}\left(\mathbbm{1}+\nu_{2}\sigma_{z}\right)$, and $\mathcal{R}_{1}=\frac{1}{2}\left(\mathbbm{1}-\nu_{2}\sigma_{z}\right)$, and consider the measurements $Q^{i}$ with $i\in\{1,2\}$, to be formed by the convex-mixing of $\mathcal{M}$ and $\mathcal{N}$ as
+Consider the measurements $\{Q^{i}\}$ with $i\in\{1,2\}$ for two possible permutations of outcomes, to be formed by the disjoint-convex-mixing of $\{M\}$ and $\{N\}$ as
 $Q^{1}=\{p\,\,M_{0}+(1-p)\,\,N_{0},\,\,pM_{1}+(1-p)N_{1}\},$
-$Q^{2}=\{p\,\,M_{0}+(1-p)N_{1},p\,\,M_{1}+(1-p)N_{0}\}.$
-In order to witness the incompatibility status of $(Q^{i},R)$, one can construct a (2,2,2) RAC task.
-Now, when Bob has to guess Alice’s 1st bit, he performs the measurement $Q^{i}$ which can be realized by the convex mixing of measurements $\mathcal{M}$ and $\mathcal{N}$ as follows: Bob tosses a biased coin which gives a ‘Head’ with probability $p$ and a ‘Tail’ with probability $(1-p)$.
-On obtaining ‘Head’, Bob performs measurement $\mathcal{M}$, and for ‘Tail’ he performs measurement $\mathcal{N}$.
-For the second bit, he performs measurement $\mathcal{R}$.
-The maximum average success probability $P(2,2,2)$ that can be obtained in the RAC task (considering $\nu_{0}=\nu_{1}=\nu_{2}$) is:
+$Q^{2}=\{p\,\,M_{0}+(1-p)N_{1},p\,\,M_{1}+(1-p)N_{0}\}.$ (59)
+To witness the incompatibility status of $(Q^{i},R)$, one can construct a $(2,2,2)-$RAC task.
+Now, when Bob has to guess Alice’s 1st bit, he performs the measurement $\{Q^{i}\}$ defined by (V.2), which is realized by the disjoint-convex-mixing of measurements $\{M\}$ and $\{N\}$.
+For the second bit, he performs the measurement $\{R\}$.
+The maximum average success probability $P(2,2,2)$ that can be obtained in the RAC task is:
 $P(2,2,2)$ $=$ $\frac{1}{8}\sum_{j,k=0}^{1}||Q^{i}_{j}+R_{k}||$ (60)
 $=$ $\frac{1}{4}(2+\sqrt{2\nu^{2}-2p\nu^{2}+2p^{2}\nu^{2}}).$
-It can be shown that for all possible convex-mixing the maximum average success probability is the same as Eq.(62).
-Now, to be fully incompatible w.r.t convex-mixing $P(2,2,2)>\frac{3}{4},\,\,\text{which implies}\,\,\nu^{2}(p^{2}-p+1)>\frac{1}{2}$ (by Eq.(62)).
-Since the minimum value of $(p^{2}-p+1)$ is $\frac{3}{4},$ so for $\nu>\sqrt{\frac{2}{3}}$ the measurements are fully incompatible w.r.t convex-mixing.
-In Observation 4, we have found that this kind of three qubit measurements with noise become fully incompatible w.r.t convex mixture for $\nu>\sqrt{\frac{2}{3}}$.
-So, we can conclude that when these three measurements are fully incompatible w.r.t convex mixing, quantum advantage is obtained in a suitable Random Access Code task.
+It can be shown that for all possible disjoint-convex-mixing, the maximum average success probability is the same as Eq.
+(60).
+Now, to be fully incompatible w.r.t.
+disjoint-convex-mixing $P(2,2,2)>3/4,\,\,\text{which implies}\,\,\nu^{2}(p^{2}-p+1)>\frac{1}{2}$ (by Eq.(60)).
+Since the minimum value of $(p^{2}-p+1)$ is $\frac{3}{4},$ so for $\nu>\sqrt{2/3}$ the measurements are fully incompatible w.r.t.
+disjoint-convex-mixing.
 […]
-Examples.— Consider the following three rank-one projective measurements acting on $\mathbb{C}^{3}$:
+Let us consider another example of three rank-one projective measurements acting on $\mathbb{C}^{3}$:
 $X=\{|0_{x}\rangle\langle 0_{x}|,|1_{x}\rangle\langle 1_{x}|,|2_{x}\rangle\langle 2_{x}|\},$
 $Y=\{|0_{y}\rangle\langle 0_{y}|,|1_{y}\rangle\langle 1_{y}|,|2_{y}\rangle\langle 2_{y}|\},$
-and $Z=\{|0_{z}\rangle\langle 0_{z}|,|1_{z}\rangle\langle 1_{z}|,|2_{z}\rangle\langle 2_{z}|\},$ (63)
+$Z=\{|0_{z}\rangle\langle 0_{z}|,|1_{z}\rangle\langle 1_{z}|,|2_{z}\rangle\langle 2_{z}|\},$ (61)
 where,
@@
 $|0_{z}\rangle$$\equiv$ $|0\rangle,|1_{z}\rangle\equiv|1\rangle,|2_{z}\rangle\equiv|2\rangle.$ (62)
-We take the following convex mixture of the measurements $X$ and $Y$: $A=\{p|0_{x}\rangle\langle 0_{x}|+(1-p)|0_{y}\rangle\langle 0_{y}|,p|1_{x}\rangle\langle 1_{x}|+(1-p)|1_{y}\rangle\langle 1_{y}|,p|2_{x}\rangle\langle 2_{x}|+(1-p)|2_{y}\rangle\langle 2_{y}|\}$ with $0\leq p\leq 1$.
+We take the following disjoint-convex-mixing of the measurements $X$ and $Y$: $A=\{p|0_{x}\rangle\langle 0_{x}|+(1-p)|0_{y}\rangle\langle 0_{y}|,p|1_{x}\rangle\langle 1_{x}|+(1-p)|1_{y}\rangle\langle 1_{y}|,p|2_{x}\rangle\langle 2_{x}|+(1-p)|2_{y}\rangle\langle 2_{y}|\}$ with $0\leq p\leq 1$.
 We now consider the $(2,3,3)$ RAC game involving the two measurements $A$ and $Z$.
-The maximum average success probability $P(2,3,3)$ for this RAC task is given by Carmeli et al.
-(2020),
+The maximum average success probability $P(2,3,3)$ for this RAC task is given by,
 $P(2,3,3)=\frac{1}{18}\sum_{i,j=0}^{2}||A_{i}+Z_{j}||,$ (63)
 which turns out to be greater than $\frac{2}{3}$ i.e. $P_{CB}(2,3,3)$ for all $p\in[0,1]$. Hence, the measurements $A$ and $Z$ are incompatible.
-Next, consider the following measurement (taking an arbitrary convex mixture of $X$ and $Y$) $A_{i,j,k,l,m,n}=\{p|i_{x}\rangle\langle i_{x}|+(1-p)|j_{y}\rangle\langle j_{y}|,p|k_{x}\rangle\langle k_{x}|+(1-p)|l_{y}\rangle\langle l_{y}|,p|m_{x}\rangle\langle m_{x}|+(1-p)|n_{y}\rangle\langle n_{y}|\}$ with $0\leq p\leq 1$, $i,j,k,l,m,n\in\{0,1,2\}$, $i\neq k$, $k\neq m$, $m\neq i$ and $j\neq l$, $l\neq n$, $n\neq j$.
-Note that the aforementioned measurement $A$ is nothing but $A_{0,0,1,1,2,2}$ in the present notation.
+Next, consider the following measurement (taking an arbitrary disjoint-convex-mixing of $X$ and $Y$) $A_{i,j,k,l,m,n}=\{p|i_{x}\rangle\langle i_{x}|+(1-p)|j_{y}\rangle\langle j_{y}|,p|k_{x}\rangle\langle k_{x}|+(1-p)|l_{y}\rangle\langle l_{y}|,p|m_{x}\rangle\langle m_{x}|+(1-p)|n_{y}\rangle\langle n_{y}|\}$ with $0\leq p\leq 1$, $i,j,k,l,m,n\in\{0,1,2\}$, $i\neq k$, $k\neq m$, $m\neq i$ and $j\neq l$, $l\neq n$, $n\neq j$.
+Note that the aforementioned measurement $A$ is denoted by $A_{0,0,1,1,2,2}$ following the present notation.
 Following a similar calculation, it can be shown that any such $A_{i,j,k,l,m,n}$ is incompatible with $Z$ for all $p\in[0,1]$.
 Similarly, taking an arbitrary convex combination of $X$ and $Z$ (or, $Y$ and $Z$), it can be shown that the new measurement is incompatible with $Y$ (or, $X$) for all $p\in[0,1]$.
-A generic description for the semi-device independent witness of incompatibility.— Consider a measurement assemblage of $n$ measurements: $M=\{M_{0},M_{1},\ldots,M_{n-1}\}.$ We make $m$ partitions $S=\{S_{i}\}_{i=0}^{m-1}$, $\bigcup_{i}S_{i}=M,\,\text{and}\,S_{i}\bigcap S_{j}=\phi,\,\,\forall\,\,i,j\,\,\text{with}\,\,i\neq j.$ Our purpose is to operationally witness the incompatibility of $m$ measurements produced from the convex mixing of $n$ measurements from each of the $m$ partitions.
-We can construct a $(m,d,d)$ RAC game where Alice has an $m$ dit input message, viz., $x=(x_{0},x_{1},\ldots,x_{m-1})$.
-Depending upon the message, she encodes it in a qudit and sends it to Bob, who on the other hand, gets input $y\in\{0,1,\ldots,m-1\}$ and accordingly, he has to predict the value of the corresponding bit $x_{y}$.
-He performs the measurement which is obtained by the convex mixtures of the measurements from the partition $S_{y}$ and declares the outcome of the measurement.
-Now, if the success probability $P(m,d,d)$ is greater than $P_{CB}(m,d,d)$ Saha et al.
-(2023) for all possible convex mixtures, we can operationally detect that the $m$ partitioned measurements are incompatible under convex mixing.
-V Conclusions
-Measurement incompatibility is the spectacle in quantum theory that a set of measurements cannot be performed jointly on arbitrary systems.
+A generic description for the semi-device independent witness of incompatibility.— Consider a measurement assemblage of $n$ measurements having $\overline{d}$ outcomes, =$\mathcal{M}=\{M_{z|x}\}_{z,x}$ where $x\in[n],z\in[\overline{d}]$, and the measurements act on $\mathbbm{C}^{d}$.
+We make $k$ partitions: $S=\{S_{i}\}_{i=0}^{k-1}$, $\bigcup_{i}S_{i}=\mathcal{M},\,\text{and}\,S_{i}\bigcap S_{j}=\phi,\,\,\forall\,\,i,j\,\,\text{with}\,\,i\neq j.$ Our purpose is to operationally witness the incompatibility of $k$ measurements produced from the disjoint-convex-mixing of $n$ measurements from each of the $k$ partitions.
+We can construct a $(k,\overline{d},d)-$RAC where Alice has an $k$ dit input message, viz., $x=(x_{0},x_{1},\ldots,x_{k-1})$.
+Depending upon the message, she encodes it in a qudit and sends it to Bob, who on the other hand, gets input $y\in[k]$ and accordingly, he has to predict the value of the corresponding bit $x_{y}$.
+He performs the measurement, which is obtained by the disjoint-convex-mixing of the measurements from the partition $S_{y}$ and declares the outcome of the measurement.
+Now, if the success probability $P(k,\overline{d},d)$ is greater than $P_{CB}(k,\overline{d},d)$ Saha et al.
+(2023) for all possible disjoint-convex-mixing and permutations of measurement outcomes, we can operationally witness $k-$incompatibility under disjoint-convex-mixing.
+VI Conclusions
+Measurement incompatibility is a feature in quantum theory that a set of measurements cannot be performed jointly on arbitrary systems Gühne et al.
+(2023).
 It is one of the fundamental ingredients for non-classical correlations and merits of quantum information science.
-Incompatibility offers a complex structure with different layers as the number of measurements and the dimension on which the measurements act increases.
-Thus, understanding the different levels of incompatibility with respect to elementary classical operations is of paramount importance both from the foundational and practical perspectives.
-In this work, we have considered the two most general classical operations viz., coarse-graining of different outcomes of measurements, and convex-mixing of different measurements.
-Through our present analysis we have investigated the different levels of incompatibility arising under the above classical operations.
-Since environmental effects are ubiquitous in practical scenarios, here we have investigated the tolerance thresholds for maintaining measurement incompatibility against noise under classical operations.
+Incompatibility offers a complex structure with different layers as the number of measurements and the dimension on which the measurements act increases Uola et al.
+(2021); Egelhaaf et al.
+(2024); Rahimi-Keshari et al.
+(2021).
+Thus, understanding the different levels of incompatibility with respect to elementary classical operations is of paramount importance from both the foundational and practical perspectives.
+In this work, we have considered the two most general classical operations viz., coarse-graining of different outcomes of measurements, and disjoint-convex-mixing of different measurements.
+Through our present analysis, we have investigated the different levels of incompatibility arising under the above classical operations.
+Since environmental effects are ubiquitous in practical scenarios, the tolerance thresholds for maintaining measurement incompatibility against noise under classical operations are investigated here.
 Furthermore, we have developed a method to operationally witness different levels of measurement incompatibility in the device-independent framework involving Bell-type experiments, and also in the semi-device-independent framework involving prepare-and-measure experiments.
@@
 Our results are particularly useful for the purpose of comparing measurements in terms of their degree of incompatibility.
-For example, measurements that remain fully incompatible with respect to coarse-graining of outcomes (or convex-mixing of measurements) show stronger incompatibility compared to those that are not fully incompatible, thus facilitating legitimate choice of measurements in an information-processing task where a high degree of incompatibility is required.
-Our present study motivates future work on several open directions.
-The condition for full incompatibility for two projective measurements with respect to coarse-graining of outcomes can be extended for more number of projective measurements.
-Similarly, the criterion for full incompatibility of projective measurements with respect to convex mixing could be generalized for any set of measurements.
-Further investigation would be required to see if device-independent witnesses could be formulated for more than three outcome measurements with respect to coarse-graining and more than three measurements with repect to convex mixing.
-It will be interesting to look for examples where the full incompatibility with respect to coarse-graining can be inferred in a device-independent way from a single experimental statistics.
-VI Acknowledgements
+For example, measurements that remain fully incompatible with respect to coarse-graining of outcomes (or disjoint-convex-mixing of measurements) show stronger incompatibility compared to those that are not fully incompatible, thus facilitating the legitimate choice of measurements in an information-processing task where a high degree of incompatibility is required.
+Our work allows for experimental implementation and one can infer more about these layers of incompatibility by employing these operations in the statistics of previously conducted experiments Zhang et al.
+(2024).
+Further, our formalism can be applied to predict the threshold values of system parameters enabling observations of incompatibility layers in future experiments.
+Our present study motivates future work in several open directions.
+The condition for full incompatibility for two projective measurements with respect to the coarse-graining of outcomes can be extended for more projective measurements.
+Similarly, the criterion for full incompatibility of projective measurements with respect to disjoint-convex-mixing could be generalized for any set of measurements.
+It will be interesting to look for examples where the full incompatibility with respect to coarse-graining can be inferred in a device-independent way from single experimental statistics.
+Acknowledgements
 […]